### v1

# DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning

Abstract.
The effectiveness of long-context modeling is important for Large Language Models (LLMs) in various applications. Despite their potential, LLMs’ efficacy in processing long context does not consistently meet expectations, posing significant challenges for efficient management of prolonged sequences in training. This difficulty is compounded by the scarcity of comprehensive and diverse training datasets suitable for long sequences, which stems from inherent length biases across different data sources, and the logistical complexities associated with massive data management for training in extended contexts. In this work, we introduce DataSculpt, a data construction framework designed to strategically augment the data architecture for extended-context training. Our thorough evaluations demonstrate DataSculpt’s remarkable capacity to boost long-context training performance, achieving improvements including an 18.09% increase in retrieval augmentation, 21.23% in summarization, 21.27% in reading comprehension, and a 3.81% rise in code completion, all while preserving the models’ overall proficiency with a 4.88% improvement.
PVLDB Reference Format:
Keer Lu, Zheng Liang, Xiaonan Nie, Da Pan, Shusen Zhang, Keshi Zhao, Weipeng Chen, Zenan Zhou, Guosheng Dong, Wentao Zhang, Bin Cui. PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX Corresponding Author.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and other artifacts will be released after the paper is accepted.
1. Introduction
Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) (Brown et al., 2020; Achiam et al., 2023; Lewkowycz et al., 2022), particularly for Natural Language Processing (NLP) tasks, reshaping AI research and applications with unparalleled success across benchmarks and real-world scenarios (Dwivedi et al., 2021). Long context sequence enables LLMs to perform tasks significantly beyond existing paradigm (Fu et al., 2024), making it a pivotal area of research that has garnered substantial attention in recent years (Vaswani et al., 2017). However, LLMs’ efficacy in processing long context does not consistently meet the expected standards. Furthermore, despite claims of its feasibility, simply extending the training context window of LLMs does not guarantee an improvement in the long context modeling capabilities (Li et al., 2024). Hence, there is a compelling call for innovative strategies and methods to effectively cultivate the long context capabilities of LLMs.
Data plays a fundamental role during the training phases of LLMs (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023). While the significance of model architectures is widely acknowledged, data-focused approaches hold substantial importance (Wang et al., 2023). The strategic management of data, especially in constructing an appropriately tailored training dataset, is critical for synergizing model performance and streamlining training processes across both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017). Efforts in data management have been directed towards exploring ways to enhance model capabilities for processing long contexts. ICLM (Shi et al., 2023) situates related documents within the same context, thereby exposing LLMs to extensive relevant contexts and supplying training indicators that transcend document boundaries. Similarly, SPLICE (Staniszewski et al., 2023) utilized retrieval methodologies such as BM25 (Robertson et al., 2009) and Contriever-MSMARCO (Izacard et al., 2021) to execute structured packing within the training context, with a focus on code-related data. However, given that each document in the training corpus is accessed only once, the strategy employed in the previous approaches yields a notable quantity of isolated documents. As a result, during the later stages, this leads to the concatenation of unvisited documents with relatively low semantic correlation. Simultaneously, considering the current state of extended context sequences used for training, it leads to three key challenges:
Figure 1. Distribution of token counts for individual documents across different domains. Domains are arranged in descending order from top to bottom based on the proportion of documents with token numbers exceeding 64K. Documents within the domains of books and academic papers predominantly feature medium to long lengths (L $\geq$ 16K), in contrast to other fields, especially web-sourced data, where document lengths are almost exclusively below 4K.
C1. Insufficient Domain-Balanced Long Training Corpora. The challenge lies in effectively incorporating a wide range of domains into extended-context LLM training. Achieving a balanced domain mixture is imperative, since each domain encompasses distinct knowledge. However, the distribution of text lengths varies substantially across various domains, as depicted in Figure 1. While some domains offer data lengths suitable for long-context training, other domains, particularly web-sourced data where less than 1% of documents exceed 4K tokens, necessitate the concatenation of multiple documents to achieve adequate training contexts. This may lead to increased cross-attention among different documents within sequences, which is an undesirable learning signal.
C2. Absence of Criteria for Long Context Construction. While existing work has demonstrated the positive impact of structured data on amplifying language models’ long-context capabilities (Levine et al., 2021; Chan et al., 2022; de Vries, 2023a; Devlin et al., 2018), these efforts remain preliminary. Research dedicated to the organization and quality assessment of long-sequence training data for LLMs tends to focus on a narrow set of factors, lacking a systematic framework. Consequently, there is a discernible need for a detailed and cohesive framework that systematically organizes training data for applications requiring extensive context.
C3. Efficiency in Massive Data Management. Training LLMs with extended contexts from scratch is notably resource-intensive, requiring considerable time and labor investments (Chen et al., 2024a), which has redirected emphasis towards post-training optimizations as a more feasible strategy. Furthermore, organizing extensive pretraining datasets into long-context sequences poses substantial logistical hurdles (To et al., 2018), further compounding the complexity of developing LLMs efficiently. Consequently, devising and implementing strategies for efficient data scheduling and management is essential to enhance the capabilities of LLMs in a sustainable and cost-effective manner (Polyzotis et al., 2017), balancing the need for extensive context processing with the practical limitations of resource availability.
To address these issues, we introduce DataSculpt, a novel framework designed to reconfigure the organization of extended-context training data into a multi-objective combinatorial optimization problem, as illustrated in Figure 2. DataSculpt meticulously balances several critical objectives throughout the data organization process, including relevance, homogeneity, integrity, as well as computational efficiency, adopting a coarse-to-fine methodology to optimize the architecture of training data. We conducted post-training with 15 billion tokens on a 7B model across varying sequence lengths of 16K, 32K, and 64K, and experimental results demonstrate that DataSculpt improves performance in an array of tasks, most significantly in general understanding (+4.88%), retrieval augmentation (+18.09%), summarization (+21.23%), reading comprehension (+21.27%) and code completion (+3.81%). The finding reveals that DataSculpt can markedly synergize the model’s performance in long-context utilization during the post-train phase, even with a constrained number of training tokens.
Our contributions can be summarized as follows:
  • We highlight the challenge of compiling an appropriate long-context training corpus with a balanced domain mixture, crucial for enhancing the capabilities of LLMs in processing extensive text sequences.
  • To create suitable long context training data, we propose an efficient framework DataSculpt, which employs a multi-objective partition allocation strategy for the construction of extended context sequences.
  • Through quantitative analysis and t-SNE visualization of clustering outcomes at the intermediary stage of DataSculpt, we validate DataSculpt’s robustness for multi-objective data concatenation within extended sequences.
  • We engaged in continual pretraining using DataSculpt with various context lengths on a 7B model, and subsequently assessed its performance across 17 distinct tasks, illuminating the profound impact of our proposed method.
2. Related Work
2.1. Training Long-Context Language Models
Training Large Language Models (LLMs) with expanded context windows facilitates the execution of tasks that surpass the capabilities defined by the existing paradigm, such as multi-document question answering (Caciularu et al., 2023), long document summarization (Bai et al., 2023), and enhanced retrieval functionalities (Liu et al., 2024). Nevertheless, the endeavor of training LLMs with extended contexts is notably time-consuming and labor-intensive. Furthermore, documents with long sequences are conspicuously scarce in the training corpus (de Vries, 2023a). For instance, less than 5% of documents within CommonCrawl exceed 2k tokens in length (Cafarella and Halevy, 2011), which hinders LLMs’ exposure to complex narratives. Despite these challenges, significant strides have been made toward enhancing LLMs’ capabilities for long context processing.
Architecture-Based Strategies.  Numerous studies explored the variants of Transformer architecture by modifying its attention mechanism, which includes introducing recurrence (Dai et al., 2019), transforming attention into less computationally demanding forms (Beltagy et al., 2020; Zaheer et al., 2020), and employing low-rank approximations (Wang et al., 2020; Peng et al., 2021). Furthermore, LONGLLAMAs (Tworkowski et al., 2024) presented the Focused Transformer (FOT) featuring a novel training objective that enhances the (key, value) structure, which facilitates the expansion of context length and enables efficient scaling through k-nearest neighbors lookup. These approaches achieve more efficient exact attention through the development of an IO-optimized CUDA kernel specifically designed for this purpose.
Data-Centric Approaches.  The effectiveness of long-context language models relies equally upon the innovative employment of data-centric strategies throughout their training processes. The construction of long-context data involves more than a mere concatenation of short documents, which can lead to cross-document attention within a sequence. In these circumstances, formulating high-quality or domain-specific long sequences holds significant research value (de Vries, 2023b). ICLM (Shi et al., 2023) and SPLICE (Staniszewski et al., 2023) combine similar training instances through a sophisticated retrieval process, which emphasizes the critical role of high-quality long-context training data in enhancing the context expansion capabilities of LLMs without necessitating extensive data resources. Additionally, Krell et al. (Krell et al., 2021) deliberate on the challenges associated with the conventional concat-and-chunk approach, proposing an approximate bin-packing algorithm as a potential solution. As an improvement, dataset decomposition (Pouransari et al., 2024) and Best-fit Packing (Ding et al., 2024) both propose the sequence length aware optimization technique to minimize truncations, thereby facilitating more efficient pretraining on long sequences and achieving scalability relative to the size of dataset. Moreover, ProLong (Chen et al., 2024a) introduces an evaluation framework for assessing the quality of long-context data, further enriching the field’s methodological toolkit. Nonetheless, these methods have not performed a systematic framework tailored for data construction in long context training. Our work conceptualizes data management for extended-context training as a multi-objective combinatorial optimization problem, and provide a comprehensive framework for long sequence data organization.
2.2. Multi-Objective Combinatorial Optimization (MOCO)
Multi-objective combinatorial optimization (MOCO) problems are characterized by the need to optimize multiple objectives simultaneously, which are often in conflict with one another (Ulungu and Teghem, 1994). The complexity of MOCO challenges lies in the fact that improving the outcome of one objective often leads to the deterioration of another, making it necessary to find a balance or trade-off between them.
MOCO-Augmented LLMs.  Existing works have integrated MOCO problems with relevant concerns in the LLMs field. Wang et al. (Wang et al., 2024) employed the MOCO approach within the Reinforcement Learning from Human Feedback (RLHF) framework for LLMs, where the reward models aggregate multiple reward objectives with coefficients into a scalar score. Akiba et al. (Akiba et al., 2024) viewed model merging as a MOCO challenge, and introduced an innovative approach that leverages evolutionary algorithms to facilitate the automated development of robust foundation models. Furthermore, integration of MOCO with data-centric LLMs have been explored for training data selection (Du et al., 2023; Xu et al., 2023; Chen et al., 2024a), which applies heuristic techniques to assign weights to different data evaluation metrics and facilitates a phased resolution approach. Considering the extensive scale of training dataset, data construction for long context is treated as a MOCO challenge, for which the coarse-to-fine algorithm is identified as a suitable resolution strategy.
3. Method
We propose an innovative approach to construct concatenated sequences with multiple attributes for the enhancement of LLMs long-context utilization.
3.1. Multi-Objective Partition Allocation
Figure 2. Illustration of DataSculpt. In the data preprocessing stage, documents are initially divided into chunks with $\leq$ context length $L$, which are subsequently transformed into vector embeddings. During the semantic clustering phase, we implement a variant of the ISODATA algorithm, which is augmented with the FAISS vector searching library, to aggregate documents based on semantic similarity. Following this, a greedy semantic-driven largest-fit algorithm is utilized to arrange these clusters into sequences that are multi-objective optimally configured for extended-context training.
We consider the problem as a multi-objective partition allocation problem with soft constraints in a massive data scale, which is essentially an instance of the multi-objective combinatorial optimization (MOCO) problem (Lin et al., 2022). This can be formulated as
$\max_{x\in\chi}\mathbf{f}\left(x\right)$ $:=\left(f_{1}\left(x\right),f_{2}\left(x\right),...,f_{m}\left(x\right)\right)$
(1) s.t. $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k},k=1,...,n$
where $x$ denotes a solution in the decision space $\chi\subset\mathbb{R}^{n}$, $\mathbf{f}:\chi\to\mathbb{R}^{m}$ represents an $m$-dimensional vector-valued objective function, and $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k}\,\text{where}\,k=1,...,n$ stands for a series of soft constraints. Importantly, the evaluation process is computationally expensive for each individual objective $f_{i}\left(x\right),i=1,...,m$.
Setup.  Suppose that we have a training corpus $\mathcal{D}$ and training sequences $\mathcal{W}$ pending sampling, the context window for each is limited in length $L$. The multi-objective partition allocation problem encompasses the following optimization objectives, the ultimate goal of which is to efficiently amplify the homogeneity of concatenated documents within the identical attention sequence:
  • G1: Relevance. One fundamental goal is to maximize the holistic semantic coherence among documents contained in the same context window, with the observation that indicates a reduction in the perplexity of language models with increasing relevance (Shi et al., 2023).
  • G2: Homogeneity. Another objective to be pursued is diminishing cross-document attention during the model’s training phase, which can be accomplished by mitigating the quantity of documents utilized for concatenation within each sequence.
  • G3: Integrity. Eliminating unnecessary truncation when concatenating documents for each training sequence is essential for preserving the integrity of individual documents (Ding et al., 2024), ensuring that the contextual richness and informational completeness are maintained to the greatest extent possible.
  • G4: Compute Efficiency. In the process of managing vast training datasets for LLMs, efficiency emerges as a critical consideration. When aligning long-context training data with multi-objective goals, a balance among accuracy, processing speed, and space consumption is vital for enhancing LLM performance while optimizing resource efficiency.
Soft constraints mandate that the sum of document length within each individual context window can not surpass the respective capacity, with the aim of maximizing document integrity and minimizing truncated token proportion within the same document (G3). Thus, Equation 1 can be written as
(2) $\max_{x\in\chi}\mathbf{f}\left(x\right):=\alpha f_{1}\left(x\right)+\beta f_{2}\left(x\right)+\lambda p\left(x\right)$
where soft constraints are transformed into penalty terms, while $\alpha,\beta,\lambda$ are non-negative hyper-parameters that control the impact of objective and penalty functions.
The respective objective functions are
(3) $f_{1}\left(x\right)$ $=\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}\sum_{\begin{subarray}{c}d_{j}\in\mathcal{D}\\ j\neq i\end{subarray}}x_{d_{i}w}x_{d_{j}w}\cos\left(E(d_{i}),E(d_{j})\right)$
$f_{2}\left(x\right)$ $=\sum_{w\in\mathcal{W}}\frac{L}{\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}$
where $f_{1}\left(x\right)$ aims to ensure the overall similarity of concatenated documents (G1), while $f_{2}\left(x\right)$ seeks to guarantee the homogeneity within a training sequence (G2). $E(d_{i})$ denotes the embedding of document $d_{i}\in\mathcal{D}$ employed embedding model (Chen et al., 2024b) and cosine similarity is used to determine the similarity between any pair of documents, while $x_{d_{i}w}$ equals 1 if $d_{i}\in w$ and 0 otherwise. $L$ is used to divide the objective function $f_{2}\left(x\right)$ to transform the minimization problem into a maximization problem for resolution.
Moreover, the penalty term $p(x)$ can be written as
(4) $p(x)=\begin{cases}0,\text{if }\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}\leq L&\\ \frac{L}{\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}-1,\text{otherwise}&\end{cases}$
which serves as a penalty for the truncation of concatenated documents that exceed context window capacity, thereby causing document incompleteness.
Solving MOCO problems exactly is NP hard, but coarse-to-fine algorithms are recognized for offering an efficient approximate solution (G4). However, the challenge lies in effectively reconciling multiple objects within the constraints of computational resources and the vast scale of the training corpus $\mathcal{D}$ (Zhao et al., 2005). We denote $|\mathcal{D}|$ as the number of documents contained in $\mathcal{D}$. The quantity of documents $k$ chosen for concatenation within a specific context is notably smaller than the total document count in $\mathcal{D}\,(k<<|\mathcal{D}|)$. Consequently, despite conducting score-based sorting on the remaining $|\mathcal{D}|-k$ documents during each iteration, they remain unselected, thereby consuming a significant expenditure of time and computational resources. Therefore, we initially perform semantic clustering on the entire training corpus $\mathcal{D}$ at a coarser granularity, followed by the implementation of a greedy algorithm to optimize the trade-off between efficiency and precision during the document organization process.
3.2. Phase 0: Data Preprocessing
Given the training context window length $L$ of the model, we tokenize raw documents in the training corpus $\mathcal{D}$ and divide them into segments, each comprising no more than $L$ tokens, which is dictated by the context length constraints. Subsequently, we utilize the embedding model to embed each segmented document $d_{i}$ to get $E\left(d_{i}\right)$.
3.3. Phase 1: Semantic Clustering
Algorithm 1 Variant of ISODATA
Input: Training corpus $\mathcal{D}$, Cluster number $N_{c}$ gained from Equation 6, Hyperparameters: distance threshold $\delta$, alteration threshold $\epsilon$, maximum iterations $T$
Parameter: Unclustered data $\mathcal{U}$, Alteration distance $dst$
Output: Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$
Define $\mu(c)$: center vector of cluster $c$
1:  Sample $N_{c}$ centroids in $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ randomly from $\mathcal{D}$, $\mu(c_{j})\leftarrow E(d_{j})\,\text{for all }j=1,...,N_{c}$
[…]
3:  for $t=1,2,\ldots,T$ do
4:     for each documnet $d_{i}$ in $\mathcal{D}$ do
5:        if $\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))>\delta$ or $t==T$ then
6:           Add $d_{i}$ to $c_{j^{*}}=\arg\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))$
7:        else
8:           $\mathcal{U}.append(d_{i})$
9:        end if
10:     end for
11:     for each $d_{u}$ in $\mathcal{U}$ do
12:        Add new cluster $c_{u}$, $\mu(c_{u})\leftarrow E(d_{u})$, $N_{c}\leftarrow N_{c}+1$
13:     end for
14:     for each cluster $c_{j}$ do
15:        $dst\leftarrow dst+||\mu(c_{j})-\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})||$
16:        Recalculate centroid $\mu(c_{j})\leftarrow\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})$
17:     end for
18:     for each pair of clusters $c_{i}$ and $c_{j}$ do
19:        if $cos(\mu(c_{i}),\mu(c_{j}))>\delta$ then
20:           Merge cluster $c_{i}$ and cluster $c_{j}$, $N_{c}\leftarrow N_{c}-1$
21:        end if
22:     end for
23:     if $dst<\epsilon$ then
24:        Break
25:     end if
26:     $dst\leftarrow 0$
27:  end for
28:  return  Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ where $\cup_{j=1}^{N_{c}}c_{j}=\mathcal{D}$
In the initial phase, the documents undergo a process of semantic clustering. To estimate the initial cluster number, we derive inspiration from the construction of a distance matrix by quantifying semantic distances between documents within the training corpus $\mathcal{D}$. We first divide $\mathcal{D}$ into $n$ subsets $D_{1},D_{2},...,D_{n}$, where each subset $D_{j}=\{d_{1},d_{2},...,d_{k}\}\subset\mathcal{D}$ and $\bigcup_{j=1}^{n}D_{j}=\mathcal{D}$. For each subset $D_{j}$, we compute the density $\rho_{j}$ of its distance matrix, which demonstrates the average dissimilarity score in relation to the entire subset.
(5) $\rho_{j}=\frac{2}{n(n-1)}\sum_{d_{i}\in D_{j}}\sum_{\begin{subarray}{c}d_{k}\in D_{j}\\ k\neq i\end{subarray}}\cos\left(E(d_{i}),E(d_{k})\right)$
Ultimately, the approximate number of clusters $N_{c}$ for semantic clustering is ascertained to be equivalent to the total number of documents within the training corpus multiplied by the average score of dissimilarity among subsets $D_{1},D_{2},...,D_{n}$, denoted as
(6) $N_{c}=\left\lfloor|\mathcal{D}|*\bar{\rho}\right\rfloor\\ =\left\lfloor|\mathcal{D}|*\frac{\sum_{D_{j}\subset\mathcal{D}}\rho_{j}}{n}\right\rfloor$
We have implemented modifications to the ISODATA clustering algorithm (Memarsadeghi et al., 2007), tailoring the algorithm’s splitting and merging phases to better align with the specific characteristics of our text corpus, as outlined in Algorithm 1. Furthermore, to enhance the efficiency of vector similarity assessments, we have integrated the FAISS index (Johnson et al., 2019), a sophisticated method for conducting high-speed pairwise similarity searches.
We first randomly select $N_{c}$ data samples from the entire training corpus $\mathcal{D}$ to serve as the initial clustering centers, which is based on the calculations from Equation 6. These sampled data points are then indexed using the FAISS vector searching library. Throughout the clustering process, for each document $d_{i}$, we employ the HNSW method (Malkov and Yashunin, 2018) from FAISS to retrieve its nearest cluster center $c_{j}$. If the semantic similarity between the current center and $d_{i}$ exceeds a predefined threshold $\delta$, then $d_{i}$ is assigned to cluster $c_{j}$. Conversely, if the similarity does not surpass $\delta$, $d_{i}$ is designated as a new cluster center. Moreover, as the iterative clustering progresses, adjustments to the cluster centers are made by calculating the mean of the document embedding vectors within each cluster, thereby establishing new cluster centroids. Additionally, if the similarity between any two clusters exceeds the threshold $\delta$, those clusters are merged. This approach not only refines the clustering process by dynamically adjusting cluster centers based on semantic coherence but also ensures that clusters remain distinct and meaningful by merging overly similar clusters, thereby enhancing the overall structure and interpretability of the clustering outcome.
Time Complexity.  The computational complexity of the original ISODATA algorithm is denoted as $\mathcal{O}(|D|TN_{c})$, where $|D|$ represents the quantity of the entire dataset, $N_{c}$ denotes the initial cluster count, and $T$ signifies the number of iterations. However, the integration of FAISS library leveraging HNSW search significantly accelerates the querying for relative vectors through the establishment of a multi-layered indexing architecture, which optimizes the process of pinpointing similar data points across expansive datasets. The construction phase of the index is typically completed within a time frame of $O(N_{c}\log N_{c})$, which then supports the rapid search for each query vector. By executing a hierarchical traversal within the index structure, the algorithm efficiently locates the nearest neighbor, showcasing a time complexity of $\mathcal{O}(\log N_{c})$. This optimization reduces the overall time complexity to $\mathcal{O}(|D|T\log N_{c})$, representing a critical step towards improving the computational efficiency of our data scheduling procedures, enabling more rapid and precise clustering outcomes (G4).
3.4. Phase 2: Greedy Policy
In the preliminary phase, documents within the training corpus $\mathcal{D}$ are clustered based on their semantic similarities at a broader level of granularity. The subsequent stage leverages a sophisticated multi-objective partition allocation algorithm, strategically applied across these preliminary clusters. The purpose of this algorithm, as depicted in Algorithm 2, is to systematically arrange documents into various context sequences for concatenation.
Algorithm 2 Greedy Semantic-Driven Largest-Fit Allocation
Input: Cluster $c_{j}\in\mathcal{C}$, Context sequence capacity $L$, Context sequence quantity $N_{w}$, Hyperparameters: impact of objective and penalty function $\alpha,\,\beta,\,\lambda$
Parameter: Available context sequence $\mathcal{A}$, Semantic relevance $f_{1}$, Residual capacity proportion $f_{2}$, Truncation penalty $p$, Concatenation score $F$
Output: Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$
Define $r(s)$: remaining capacity of context sequence $s$
Define $l(d)$: token quantity of document $d$
Define $\mu(s)$: center vector of context sequence $s$
1:  $\left\{r(s_{l})\leftarrow L\right\}_{l=1}^{N_{w}}$: Initialize empty context sequences
2:  $S_{c_{j}}$: Sort $d_{i}\in c_{j}$ by token quantity in descending order
3:  for each documnet $d_{i}$ in $S_{c_{j}}$ do
4:     $\mathcal{A}=\left\{s_{l}|r(s_{l})\geq 0\right\}_{l=1}^{N_{w}}$
5:     if $\mathcal{A}=\emptyset$ then
6:        Break
7:     end if
8:     for each context sequence $s_{a}$ in $\mathcal{A}$ do
9:        $f_{1}(d_{i},s_{a})=cos(E(d_{i}),\mu(s_{a}))$
10:        $f_{2}(d_{i},s_{a})=r(s_{a})/L$
11:        $p(d_{i},s_{a})=\begin{cases}0,\text{if }l(d_{i})\leq r(s_{a})&\\ L/(L+l(d_{i})-r(s_{a})),\text{otherwise}&\end{cases}$
12:        $F(d_{i},s_{a})=\alpha f_{1}(d_{i},s_{a})+\beta f_{2}(d_{i},s_{a})+\lambda p(d_{i},s_{a})$
13:        Add $d_{i}$ to $s_{a^{*}}$ where $s_{a^{*}}=\arg\max_{s_{a}\in\mathcal{S}}F(d_{i},s_{a})$
14:     end for
15:     $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$
16:     $\mu(s_{a^{*}})\leftarrow((|s_{a^{*}}|-1)\mu(s_{a^{*}})+E(d_{i}))/|s_{a^{*}}|$
17:  end for
18:  return  Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$
As delineated in Section 3.1, our objective encompasses augmenting the relevance (G1), homogeneity (G2), and integrity (G3) of the textual content encapsulated within each context window. This enhancement is realized through a strategic reduction in the quantity of documents concatenated into a single context sequence, amplifying the semantic congruence among these documents, and eliminating unnecessary truncations.
Initially, documents within each previously formed cluster are organized in descending order by their token count. Following this organization, for each document $d_{i}$ pending allocation, the process seeks to integrate it into an existing sequence $s_{a}$ with content length that does not exceed the maximum context window capacity $L$. The sequence $s_{a}$ is selected based on achieving the highest composite score $F(d_{i},s_{a})$, which reflects a weighted consideration of semantic similarity (G1), available residual capacity (G2), and the minimization of truncation penalties (G3). Through this approach, we aim to ensure that the allocation of documents to sequences is not only computing efficient but also optimally maintains semantic consistency, maximizes capacity utilization, and preserves the integrity of individual documents within each training sequence. The strategy is pivotal in enhancing the overall effectiveness of context sequences for LLMs, by fostering a training environment that is both informationally rich and contextually coherent.
4. Experiments and Results
In this section, we describe details of our analytical study of training corpuses (Section 4.1), experimental setup (Section 4.2), the baseline methods we use for comparison (Section 4.3), and experimental results (Section 4.4).
4.1. Analytical Study on Data Corpus
The data employed for training originates from a variety of domains, including the web, books, academic papers, code repositories, etc.. The composition of these sources and their respective weight proportions are illustrated in Figure 3.
Figure 3. Domain proportion of the training corpus. We have maintained the data distribution ratios across different sources from the previous pretraining stage.
To provide further insight of our intuition, we undertook a comprehensive analysis, examining both the token quantities within individual documents among diverse domains and assessing the effectiveness of semantic clustering across different methodologies.
4.1.1. Length Analysis
We conducted a statistical analysis of the token number characteristics across various domains, as illustrated in Table 1. We also categorize the tokenized documents into different length intervals based on token quantities across each domain. This process enables us to derive the distribution of token lengths for individual documents situated within specific domains. The proportional distributions across domains are presented as a normalized stacked horizontal bar chart in Figure 1, which provides a hierarchical view of the domains according to their prevalence of longer texts. It reveals that data originating from the book and academic paper domains predominantly consist of long documents, with more than half of the data exceeding sequence lengths of 64K tokens. In contrast, academic papers tend to feature medium-length documents (ranging from 4K to 32K tokens), while web-sourced data are characterized by a preponderance of short documents, almost exclusively comprising texts shorter than 4K tokens. Consequently, it is evident that web data does not inherently lend itself to extending model training windows due to its shorter length, necessitating the concatenation of multiple individual documents to form longer sequences suitable for training.
Domain Book (Chinese) Book (English) Paper (English) Code Paper (Chinese) Web (English) Web (Chinese) Others
Max Len 24,203,838 8,241,838 10,721,906 1,000,000 1,471,297 1,267,049 80,688 21,839
Min Len 667 322 127 173 696 289 132 23
Avg Len 176,156 216,969 20,555 2,578 6,216 886 358 224
Table 1. Token statistics among domains. We conducted an analysis of the maximum, minimum, and average token counts from different domains following tokenization. Data sourced from the web tends to have a shorter average length, necessitating the concatenation of multiple documents to achieve extended sequences suitable for long-context training.
The attribute that individual documents originating from books and academic papers tend to have a longer average length significantly increases their propensity to occupy the entirety of a context sequence when training models under expanded contexts, such as 32K or 64K tokens. Conversely, web-derived data predominantly exhibits a shorter average document length compared to other domains, necessitating the concatenation of multiple documents to form sequences of adequate length during the training process of extended sequence lengths.
This observation underscores the intrinsic variability in document length across different domains. However, a balanced domain mixture is important for extend context training, where experiments conducted by Fu et al. (Fu et al., 2024) demonstrated the importance of upsampling longer sequences while preserving the domain mixture within the pretraining corpora for effective context scaling. Therefore, it is necessary to adopt flexible data organization strategies tailored to accommodate the varying length distributions across domains. Specifically, for domains characterized by longer documents, such as books and academic papers, leveraging the natural length of these documents can optimize the utilization of context windows during training. On the other hand, for web-sourced data, where a majority of the individual documents are insufficient to span the entire context sequence, a strategy involving the concatenation of multiple documents to construct longer sequences becomes imperative. This insight holds substantial significance for our work. The approach ensures that the training process effectively captures the broader context as well as takes into account the characteristics of different data sources, thereby enhancing the model’s overall ability to understand and generate coherent text across varying lengths and domains.
Considering the distribution of document lengths and proportion of each domain, we have strategically organized the web-produced English data within our training corpus, which constitutes 30.5% of the total training data.
4.1.2. Semantic Clustering Analysis
Method Cluster Number Stats of Docs Number in Clusters
Max Min Mean Median Less than 100 Single Doc
ICLM 323,428,858 26,630 1 3.458 1 322,926,996 173,231,427
DataSculpt 611,981 11,747 1 1,827 763 1,479 29
Table 2. Document statistics among clusters. The clusters formed through ICLM’s greedy traversal of an incomplete graph demonstrate a significant bias in the distribution of document quantities, with 99.8% of clusters containing fewer than 100 documents and 53.56% comprising only a single document. In contrast, clusters generated by DataSculpt’s semantic clustering approach display a more balanced distribution of document numbers.
(a) T-SNE on clusters built from Algorithm 1.
(b) T-SNE on clusters built from graph traversal in ICLM.
Figure 4. Visualization using t-SNE on document embeddings sampled from the web derived (English) training corpus, with each cluster denoted by a distinct color. Left (4(a)) presents the semantic clustering results generated by the methodology outlined in Algorithm 1, and right (4(b)) illustrates clusters where nodes are united by a shared traversal path, as identified by the greedy graph traversal algorithm discussed in ICLM (Shi et al., 2023).
In evaluating the effectiveness of semantic clustering, our analysis encompasses two key dimensions: the statistical features related to the document quantities within each cluster, and the application of t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction to project and analyze clusters. These clusters were derived from two distinct approaches: the greedy traversal of incomplete graphs as described in ICLM (Shi et al., 2023), for which we refer to the official implementation  [^1] https://github.com/swj0419/in-context-pretraining, and the semantic clustering achieved through an ISODATA variant (Algorithm 1) implemented by our DataSculpt.
Numerical Narratives.  Initially, we performed a quantitative analysis of the number of clusters generated by two distinct methodologies, as detailed in Table 2. It revealed that the clusters produced through greedy graph traversal in ICLM totaled 323,428,858, markedly surpassing the 611,981 clusters formed via semantic clustering with Algorithm 1, which suggests a more fragmented clustering approach by the former method. Moreover, clusters initiated at the commencement of the traversal process encompassed a substantially larger number of documents compared to those established later, leading to a pronounced imbalance in document distribution. Specifically, a significant skew was observed in the distribution of clusters from ICLM, with 53.56% comprising isolated single-document clusters. The dominance of such minimal-scale clusters hinders the ability to adequately fill the context window during the concatenation process for long sequences, thereby inadequately mitigating cross attention in training despite the potential for high semantic similarity among documents. Conversely, DataSculpt exhibited a more equitable distribution of cluster sizes throughout the semantic clustering process, enabling the majority of clusters containing semantically akin documents to extend to the length required for a long context sequence. This strategy efficiently capitalizes on the advantages of clustering similar documents, which significantly enhances the model’s training efficiency by promoting semantic coherence within clusters.
T-SNE Visualization Journeys.  Data visualization is crucial for helping decision-making during massive data management (Qin et al., 2019). Here we also employed the dimensionality reduction technique t-SNE to visualize both the semantic clusters generated via Algorithm 1 and those identified through greedy graph traversal mechanism on an incomplete graph as described in ICLM. For the latter approach, each document was connected to its 20 nearest neighbors, facilitating the construction of the incomplete graph.
Figure 4 clearly demonstrates the effectiveness of Algorithm 1 in achieving semantic clustering, wherein the documents within each cluster display a pronounced semantic concentration. Upon applying t-SNE for dimensionality reduction and subsequent visualization, distinct clusters can be clearly delineated, showcasing notable semantic boundaries among them. This method underscores the capacity of semantic clustering to aggregate documents with high semantic similarity, enhancing the interpretability of large datasets. However, clusters obtained through the greedy graph traversal in ICLM exhibit substantial size variances, attributed mainly to the restriction of single visitation per document and absence of a strategic path truncation policy. As a result, such clusters display a wider range of semantic cohesion internally, which leads to a diminished relational coherence among documents concatenated into the same context sequence, potentially undermining the efficacy of the clustering process. The ability of Algorithm 1 to produce semantically concentrated clusters suggests a more robust framework for handling the complexities of large-scale training datasets.
4.2. Experimental Setup
We continue pretraining a Transformer decoder-only 7B model (Dong et al., 2024) similar to LLaMA framework (Touvron et al., 2023) by different sources with the same distribution ratios to the previous pretraining stage, where we structured the data derived from the web (English) domain in accordance with a previously defined multi-objective partition allocation strategy (Algorithm 2). The details of training datasets can be found in Section 4.1.
Data Processing Details.  During the data preprocessing phase, we tokenize raw documents from the training corpus (Baichuan, 2023), which collectively encompass 855 billion tokens, and then truncated documents at varying sequence lengths of 16K, 32K, and 64K. In conducting vector embedding across our training datasets, we employed the BGE-m3 model (Chen et al., 2024b) as our embedding model, encoding each document into a $1024$-dimensional vector. The encoding process necessitated approximately 9 days using a 1,044,090 GPU cluster. Following this, we constructed an HNSW index based on FAISS index (Johnson et al., 2019) to facilitate efficient similarity search for semantic clustering (Section 3.3). The greedy semantic-driven partition algorithm (Section 3.4) within each cluster was executed on a 4114 CPU Ray cluster.
Model Details.  The model comprises a Transformer decoder with 32 attention heads across 32 layers. It features a hidden layer size of 4096, coupled with a dimensionality of 11008 for the forward propagation layers, and the base of Rotary Positional Encoding (RoPE) (Su et al., 2024) is modified to accommodate longer context. We train with 16K, 32K, 64K sequence lengths using 15B tokens for the 7B base model (recall that base model has the 16K context length), applied on 112 H800 GPUs across 14 nodes. We employ global batch sizes of 224, 112, and 56 sequences per step for context lengths of 16K, 32K, and 64K, which maintains a uniform number of tokens for model parameter updates. The learning rate is set to 3e-4 with linear warm-up over the first 1,000 optimization steps and cosine decay schedule.
4.3. Baselines
We compare DataSculpt with the following baselines: (1) Random sampling is the standard method that places randomly shuffled and concatenated documents within each input context. (2) ICLM (Shi et al., 2023) utilizes a greedy graph traversal strategy to concatenate semantically similar documents within the context sequence, but only constructing an incomplete graph and restricting each document node to a single visit during traversal due to the extensive volume of the training corpus.
4.4. Results
We perform evaluations to validate the effectiveness of DataSculpt on both common capabilities and long-context utilization across various downstream tasks, including general understanding (Section 4.4.1), retrieval augmentation (Section 4.4.2), summarization (Section 4.4.3), reading comprehension (Section 4.4.4) and code completion (Section 4.4.5).
4.4.1. General Understanding
General understanding tasks require the model to comprehend the text holistically, extracting key elements and answering queries, which focuses on broad understanding over specific task execution. We evaluate the 7B natural language models with zero-shot on MMLU (Hendrycks et al., 2021b, a), TyDiQA (Clark et al., 2020), BBH (Srivastava et al., 2022) and few-shot in-context learning on TREC (Li and Roth, 2002; Hovy et al., 2001). MMLU encompasses 57 distinct tasks featuring multiple-choice questions from diverse knowledge domains such as humanities, social sciences and hard sciences. TyDiQA is a multilingual question answer dataset containing 204K pairs across 11 topologically diverse languages, which captures language phenomena absent in English-only corpora. BBH focuses on a suite of 23 challenging tasks from BIG-Bench that require multi-step reasoning, where previous language model evaluations fell short of average human-rater performance. TREC is a text classification challenge aimed at categorizing queries into predefined classes, thereby assessing a model’s ability to accurately classify questions, facilitating the determination of suitable responses or actions. Table 3 contains more details about these tasks. Each dataset includes multiple subtasks, and each subtask comes with few-shot examples.
Dataset # Tasks (Labels) # Instances Answer Type
MMLU 57 14,079 Letter options
TyDiQA 9 1,713 Span
BBH 23 920 COT and answer
TREC 6(1st)-47(2nd) 500 Letter options
Table 3. Statistics of evaluation datasets for general understanding. The evaluation tasks cover different kinds of types, including question answering and text classification.
In Table 4, we compare our models trained via DataSculpt with other baseline models in general understanding across various context window size. DataSculpt not only aligns but also slightly outperforms the baselines, particularly in areas requiring multi-hop logical reasoning tasks (+4.88% compared to random sampling, +4.38% compared to ICLM).
Context Length Method MMLU TYDIQA BBH TREC
16K random 52.37 21.45 37.87 68
ICLM 52.03 22.27 38.52 71
DataSculpt 51.92 24.75 39.7 71.5
32K random 47.09 17.27 33.81 70.5
ICLM 48.09 18.8 29.95 67.5
DataSculpt 48.36 19.53 35.35 74.5
64K random 45.26 15.28 32.9 68
ICLM 46.69 16 33.14 69.5
DataSculpt 46.27 17.55 34.54 71
Table 4. Evaluation on general understanding tasks. Despite DataSculpt targeting to long-context utilization, it achieves performance slightly exceeds that of the baselines.
4.4.2. Retrieval Augmentation
Retrieval augmentation involves LLMs to access external knowledge sources, such as databases or text corpora to retrieve pertinent information as required, effectively extending their memory and knowledge capabilities beyond what is available in their pretrained state. In assessing the performance of models on retrieval augmentation tasks, we focus on the models’ synthetic key-value retrieval performance (Liu et al., 2024) under settings of 140 and 300 key-value pairs and conduct a zero-shot evaluation using the TriviaQA dataset (Joshi et al., 2017). Further experiments are also undertaken on the “needle in a haystack” task (Kamradt, 2023; Fu et al., 2024) with a maximum of 64K tokens to gain a comprehensive understanding of the effectiveness of these models in identifying specific information within extensive contexts. TriviaQA comprises over 650K question-answer-evidence triples and 95K question-answer pairs labeled with evidence passages curated by trivia enthusiasts. The key-value retrieval task (Liu et al., 2024) has evolved into a standard diagnostic tool for exploring contextual capabilities, where models are presented with a set of key-value pairs and tasked to retrieve a value corresponding to a specific key from a set of key-value pairs. The “needle in a haystack” task is designed to evaluate the models’ capacity for in-context retrieval within extended contexts, further testing their ability to navigate and extract pertinent information from dense information sets.
Context Length Method 140 Key-Value Pairs 300 Key-Value Pairs
$p_{1}$ $p_{35}$ $p_{70}$ $p_{105}$ $p_{140}$ Avg. $p_{1}$ $p_{50}$ $p_{100}$ $p_{150}$ $p_{200}$ $p_{250}$ $p_{300}$ Avg.
16K random 97.2 59.6 20.6 16.2 8.8 40.48 - - - - - - - -
ICLM 99.8 20.2 25.6 13.6 11.6 34.16 - - - - - - - -
DataSculpt 99 77.6 46.8 36.4 12.6 64.48 - - - - - - - -
32K random 82.4 48.8 46.8 38 15 46.2 71.8 15.8 6.4 4.4 15.2 6.6 68.6 26.97
ICLM 68 53.8 56.2 52.6 47.2 55.56 75 17.7 12.3 8.6 17.5 22.3 72.4 32.26
DataSculpt 87 91 89.4 81.6 51.4 80.08 97 26.2 16 13.6 15.6 29.8 83.6 40.26
64K random 99.6 52.4 43 49.4 45.4 57.96 99.4 26.6 26.2 22.4 4.2 5.4 45.4 32.8
ICLM 92.4 74 78.8 73.6 56.8 75.12 88 51.8 48.2 31.4 16.7 22.6 56.3 45
DataSculpt 99.8 96.1 92.4 87.7 61.5 85.5 97.4 64.4 56.8 38 24.8 30.2 72.6 54.89
Table 5. Key-value retrieval performance on dictionaries of 140 ($\approx$ 11.2K tokens) and 300 ($\approx$ 24K tokens) key-value pairs. $p_{i}$ represents the scenario where the pertinent information is located at the $i$-th position. “-” indicates instances where the result is not applicable, as the tokenized context length exceeds the model’s capacity. Each position underwent evaluation across 500 examples. The model trained with DataSculpt performs much higher accuracy on hard-to-retrieve positions.
Context Length Method K-V Retrieval TriviaQA Needle.
16K random - 86.15 32.1
ICLM - 87.86 33.7
DataSculpt - 89.58 44.9
32K random 36.59 86.31 49.4
ICLM 43.91 85.48 58.2
DataSculpt 60.17 84.12 63.7
64K random 45.38 85.85 77.4
ICLM 60.06 86.02 82.8
DataSculpt 70.20 86.84 90
Table 6. Average performance on retrieval augmentation tasks. DataSculpt outperforms the baselines on almost all datasets, with the mean improvement at 19.33%.
Figure 5. “Needle in a haystack” performance comparison, where the x-axis represents the document’s length (the “haystack”), and the y-axis reflects the position of the “needle” (a short sentence) within the document, ranging from 1K to 64K tokens. The model’s performance in reciting information from the “needle” across various document lengths and positions is assessed on a scale from 1 to 10, represented through a color gradient transitioning from red (score of 1) to green (score of 10).
Results in Table 5 shows the superior performance of DataSculpt in key-value retrieval tasks across both 140 and 300 key-value pairs configurations, showcasing improvements of up to 59.09% over random sampling and 34.33% over ICLM. Since it proves challenging for a model trained with 16K sequence length to extrapolate its capabilities to 24K, models trained under such setting did not yield measurable scores in the 300 key-value pairs task. Additionally, Figure 5 illustrates the visualization outcomes for the “needle in a haystack” task, indicating that DataSculpt notably enhances the model’s ability to extrapolate beyond the training context lengths.
4.4.3. Summarization
Summarization tasks involve generation of concise and coherent abstracts from a given source text, which evaluate a model’s capacity to distill key information, maintain the original text’s intent, and present it in a brief and comprehensible manner. Compared to traditional QA and text classification tasks, which frequently rely on local context for solutions, summarization requires a broader understanding of the entire document (Bai et al., 2023). This capability is more closely associated with LLMs’ long-context capabilities. We assess the post-trained models in a zero-shot setting on GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), and employ few-shot in-context learning on SAMSum (Gliwa et al., 2019). GovReport is a large-scale dataset for long document summarization, featuring approximately 19.5K U.S. government reports paired with expert-crafted abstractive summaries. QMSum is a benchmark for query-based multi-domain meeting summarization, encompassing annotated query-summary pairs spanning 232 meetings in diverse domains including product development, academia, and committee discussions. SAMSum is specifically curated for dialogue summarization, comprising messenger-style conversations accompanied by human-annotated summaries. Additional details about these tasks can be found in Table 7. Here we utilize the ROUGE-L (Lin, 2004) metric, an N-gram based measure extensively adopted in QA and summarization tasks. Improvements in summarization tasks achieved through DataSculpt are documented in Table 8, showcasing enhancements of 21.21% over random sampling and 18.18% over ICLM.
Dataset Source Avg Len Query Metric
GovReport Government report 8,734 $\times$ Rouge-L
QMSum Meeting 10,614
SAMSum Dialogue 6,258 $\times$
Table 7. Statistics of evaluation datasets for summarization. “Source” refers to the provenance of the dataset, and “Avg Len” is calculated based on word counts. “Query” signifies whether the dataset conforms to the query-based format.
Context Length Method GovReport QMSum SAMSum
16K random 24.51 18.6 15.83
ICLM 25.5 17.43 16.67
DataSculpt 27.66 21.41 20.41
32K random 24.04 18.42 12.23
ICLM 27.71 18.72 12.85
DataSculpt 28.83 17.87 24.04
64K random 26.75 18.92 12.47
ICLM 25.92 17.82 13.65
DataSculpt 29.42 20.25 18.41
Table 8. Evaluation on summarization tasks. DataSculpt consistently surpasses baseline models across nearly all three datasets and various context lengths.
4.4.4. Reading Comprehension
Dataset Avg Len Classification Metric
NarrativeQA 18,409 Single-Doc F1
Qasper 3,619
HotPotQA 9,151 Multi-Doc
2WikiMultihopQA 4,887
Table 9. Statistics of evaluation datasets for reading comprehension. “Avg Len” is calculated based on word counts.
Context Length Method NarrativeQA Qasper
16K random 7.46 7.53
ICLM 12.34 9.6
DataSculpt 19.33 9.83
32K random 11.09 10.55
ICLM 16.47 9.63
DataSculpt 18.33 11.35
64K random 16.53 9.16
ICLM 14.07 9.97
DataSculpt 18.45 9.07
Table 10. In zero-shot evaluations on the validation subsets of NarrativeQA and Qasper, DataSculpt records improvements of 27.06% over random sampling and 13.22% over ICLM.
Context Length Method HotPotQA 2WikiMultihopQA
16K random 3.52 8.66
ICLM 8.07 9.57
DataSculpt 7.68 9.82
32K random 10.09 10.45
ICLM 8.73 10.39
DataSculpt 8.31 11.24
64K random 8 9.98
ICLM 8.47 9.51
DataSculpt 9.92 10.23
Table 11. Zero-shot F1 performance on the reading comprehension tasks of HotPotQA and 2WikiMultihopQA datasets, with an average improvement of 10.54% on HotPotQA and 6.86% on 2WikiMultihopQA.
Figure 6. Radar charts depicting the average performance metrics of models trained with varying sequence lengths on a spectrum of benchmark categories. Among all evaluated training lengths, models developed using the DataSculpt approach surpass baseline-trained models in performance metrics across a diverse range of benchmark categories.
Reading comprehension assesses the capacity of models to generate responses to queries, grounded in the context provided within a given textual passage. We evaluate the models’ ability in reading comprehension with zero-shot on NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), HotPotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). NarrativeQA consists of long stories in the form of books and movie scripts, accompanied by human-crafted questions designed to assess reading comprehension. Qasper is a dataset comprising 5,049 questions derived from 1,585 NLP papers. The questions are formulated by NLP experts based solely on paper titles and abstracts, while answers are addressed by a different group of NLP practitioners. HotpotQA involves 113,000 two-hop question-answer pairs with four key features, authored directly by native speakers given two related paragraphs. 2WikiMultihopQA is a multi-hop dataset that has comprehensive explanations from question to answer derived from Wikipedia and Wikidata, synthesized with manually designed templates to prevent shortcut-based solutions. NarrativeQA and Qasper fall into the category of single-document question-answering (QA) datasets, whereas HotPotQA and 2WikiMultihopQA are recognized as Wikipedia-based, multi-hop QA datasets. Evaluation across these datasets is conducted using the F1 metric, as outlined in Table 9. Enhancements attributed to DataSculpt in reading comprehension tasks are detailed in Table 10 and Table 11.
4.4.5. Code Completion
Code completion serves as a critical and challenging task for evaluating a model’s long context modeling ability, which facilitates user coding by predicting and completing code segments based on previous inputs and contextual understanding. For code completion tasks, we assess the models using zero-shot learning on the HumanEval dataset (Chen et al., 2021), LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023), where the latter two datasets are specifically designed for long code modeling evaluation. HumanEval contains 164 handwritten programming problems, each accompanied by a function signature, a documentation string, the main body of the code, and an average of 7.7 unit tests per challenge to assess correctness. LCC is developed by selecting single-file code snippets from GitHub based on length criteria. RepoBench-P is curated from GitHub repositories, integrating retrieval and code completion tasks into a cohesive pipeline, where the goal is to first identify and retrieve the most pertinent code snippets based on the context within a file, and then predict the most appropriate subsequent line of code based on the retrieved snippets as well as the in-file context. Further information regarding these tasks is provided in Table 12, where we applied execution-based accuracy (Pass@1) for HumanEval evaluation and Edit Sim (Levenshtein distance) for the remaining two code generation tasks.
Dataset Language Level Metric
HumanEval Python function Pass@1
LCC Python/Java/C# file Edit Sim
RepoBench-P Python/Java repository
Table 12. Statistics of datasets for code completion. “Language” refers to the programming languages included within the dataset, and “Level” denotes the task’s focus granularity.
Context Length Method HumanEval LCC RepoBench-P
16K random 10.37 65 60.24
ICLM 10.58 64.22 55.64
DataSculpt 11.74 68.28 61.54
32K random 12.2 64.52 57.85
ICLM 12.26 62.91 60.94
DataSculpt 12.68 68.58 59.21
64K random 12.8 64.01 57.06
ICLM 12.88 67.11 58.9
DataSculpt 13.01 64.37 60.04
Table 13. Performance on code completion, with an enhancement of 4.23% over random sampling and 3.46% on ICLM.
Results of zero-shot evaluation are presented in Table 13, which reveals a modest advantage of DataSculpt in synergizing average scores for models’ code capabilities. While the data was not explicitly structured for the coding domain, advancements in long-context processing achieved through DataSculpt have correspondingly manifested in enhanced coding performance.
5. Conclusion
The ability to process long contexts is crucial for Large Language Models (LLMs), making the effective training of LLMs for extended-context tasks from a data-centric perspective a worthy area of investigation. A pressing challenge lies in organizing long-sequence data for domains lacking extensive text, while ensuring diversity in domain knowledge. In response, we introduce DataSculpt, a framework for organizing long-context training data, conceptualizing the task as a multi-objective combinatorial optimization (MOCO) problem. This framework considers factors including relevance, homogeneity, integrity, and computational costs associated with data construction, employing a coarse-to-fine method for phased greedy solutions. Experimental results indicate that data organized via DataSculpt significantly enhances long-context capabilities without compromising general inference abilities, marking a substantial improvement over existing data management methods.

### v2

@@
-DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning
+DataSculpt: Crafting Data Landscapes for Long-Context LLMs through Multi-Objective Partitioning
 Abstract.
-The effectiveness of long-context modeling is important for Large Language Models (LLMs) in various applications.
-Despite their potential, LLMs’ efficacy in processing long context does not consistently meet expectations, posing significant challenges for efficient management of prolonged sequences in training.
-This difficulty is compounded by the scarcity of comprehensive and diverse training datasets suitable for long sequences, which stems from inherent length biases across different data sources, and the logistical complexities associated with massive data management for training in extended contexts.
-In this work, we introduce DataSculpt, a data construction framework designed to strategically augment the data architecture for extended-context training.
-Our thorough evaluations demonstrate DataSculpt’s remarkable capacity to boost long-context training performance, achieving improvements including an 18.09% increase in retrieval augmentation, 21.23% in summarization, 21.27% in reading comprehension, and a 3.81% rise in code completion, all while preserving the models’ overall proficiency with a 4.88% improvement.
+In recent years, foundation models, particularly large language models (LLMs), have demonstrated significant improvements across a variety of tasks.
+One of their most important features is long-context capability, which enables them to process and generate extended text while maintaining coherence, retrieving relevant information, and efficiently handling tasks with substantial amounts of text.
+The key to improving long-context performance lies in effective data organization and management strategies that integrate data from multiple domains and optimize the context window during training.
+Through extensive experimental analysis, we identified three key challenges in designing effective data management strategies that enable the model to achieve long-context capability without sacrificing performance in other tasks: (1) a shortage of long documents across multiple domains, (2) effective construction of context windows, and (3) efficient organization of large-scale datasets.
+To address these challenges, we introduce DataSculpt, a novel data management framework designed for long-context training.
+We first formulate the organization of training data as a multi-objective combinatorial optimization problem, focusing on attributes including relevance, homogeneity, integrity, and efficiency.
+Specifically, our approach utilizes a coarse-to-fine methodology to optimize training data organization both efficiently and effectively.
+We begin by clustering the data based on semantic similarity (coarse), followed by a multi-objective greedy search within each cluster to score and concatenate documents into various context windows (fine).
+Our comprehensive evaluations demonstrate that DataSculpt significantly enhances long-context training performance, resulting in improvements of $18.09\%$ in retrieval augmentation, $21.23\%$ in summarization, $21.27\%$ in reading comprehension, and a $3.81\%$ increase in code completion, while also maintaining overall model proficiency with a $4.88\%$ improvement.
 PVLDB Reference Format:
-Keer Lu, Zheng Liang, Xiaonan Nie, Da Pan, Shusen Zhang, Keshi Zhao, Weipeng Chen, Zenan Zhou, Guosheng Dong, Wentao Zhang, Bin Cui. PVLDB, 14(1): XXX-XXX, 2025.
-doi:XX.XX/XXX.XX Corresponding Author.
+Keer Lu, Xiaonan Nie, Zheng Liang, Da Pan, Shusen Zhang, Keshi Zhao, Weipeng Chen, Zenan Zhou, Guosheng Dong, Bin Cui, Wentao Zhang. PVLDB, 18(1): XXX-XXX, 2025.
+doi:XX.XX/XXX.XX Equal Contribution.
+Corresponding Author.
 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
@@
 Publication rights licensed to the VLDB Endowment.
-Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
 doi:XX.XX/XXX.XX
 PVLDB Artifact Availability:
-The source code, data, and other artifacts will be released after the paper is accepted.
+The source code, data, and/or other artifacts have been made available at https://github.com/8023looker/DataSculpt.
 1. Introduction
-Large Language Models (LLMs) have become a cornerstone in Artificial Intelligence (AI) (Brown et al., 2020; Achiam et al., 2023; Lewkowycz et al., 2022), particularly for Natural Language Processing (NLP) tasks, reshaping AI research and applications with unparalleled success across benchmarks and real-world scenarios (Dwivedi et al., 2021).
-Long context sequence enables LLMs to perform tasks significantly beyond existing paradigm (Fu et al., 2024), making it a pivotal area of research that has garnered substantial attention in recent years (Vaswani et al., 2017).
-However, LLMs’ efficacy in processing long context does not consistently meet the expected standards.
-Furthermore, despite claims of its feasibility, simply extending the training context window of LLMs does not guarantee an improvement in the long context modeling capabilities (Li et al., 2024).
-Hence, there is a compelling call for innovative strategies and methods to effectively cultivate the long context capabilities of LLMs.
-Data plays a fundamental role during the training phases of LLMs (Miao et al., 2024; Fernandez et al., 2023; Whang et al., 2023).
-While the significance of model architectures is widely acknowledged, data-focused approaches hold substantial importance (Wang et al., 2023).
-The strategic management of data, especially in constructing an appropriately tailored training dataset, is critical for synergizing model performance and streamlining training processes across both pretraining and supervised fine-tuning stages (Wang et al., 2023; Kumar et al., 2017).
-Efforts in data management have been directed towards exploring ways to enhance model capabilities for processing long contexts.
-ICLM (Shi et al., 2023) situates related documents within the same context, thereby exposing LLMs to extensive relevant contexts and supplying training indicators that transcend document boundaries.
-Similarly, SPLICE (Staniszewski et al., 2023) utilized retrieval methodologies such as BM25 (Robertson et al., 2009) and Contriever-MSMARCO (Izacard et al., 2021) to execute structured packing within the training context, with a focus on code-related data.
-However, given that each document in the training corpus is accessed only once, the strategy employed in the previous approaches yields a notable quantity of isolated documents.
-As a result, during the later stages, this leads to the concatenation of unvisited documents with relatively low semantic correlation.
-Simultaneously, considering the current state of extended context sequences used for training, it leads to three key challenges:
+In recent years, artificial intelligence has achieved remarkable milestones across a wide range of fields (Dwivedi et al., 2021; Lewkowycz et al., 2022), including computer vision (Dosovitskiy, 2020; Liu et al., 2021; Riquelme et al., 2021), natural language processing (Brown et al., 2020; Devlin et al., 2018; Radford et al., 2019; Raffel et al., 2020) and generative AI (Achiam et al., 2023).
+These breakthroughs are the result of significant progress in the development of AI systems, which is built upon three critical components: data, algorithms, and computational resources.
+Data forms the foundation, serving as the essential input from which models extract and generalize patterns (Wang et al., 2023; Kumar et al., 2017).
+The quality, quantity, and diversity of the data are pivotal in shaping a model’s ability to perform effectively across a range of tasks.
+Algorithms define how data is processed and guide the architectural design of models, enabling them to address specific problems.
+Computational resources, such as GPUs, are vital for handling the complex computations required throughout the training process, ensuring scalability and efficiency as models grow in complexity (Nie et al., 2023; Miao et al., 2024).
+Historically, advancements in AI were primarily driven by model-centric approaches, where the focus was on refining and optimizing algorithmic architectures, such as AlexNet (Krizhevsky et al., 2012), ResNet (He et al., 2016), GoogleNet (Szegedy et al., 2015) in convolutional neural networks.
+However, in recent years, there has been a paradigm shift from the model-centric approach to the data-centric approach for foundation models, such large language models (LLMs).
+This shift emphasizes the quality, organization, and management of data as the critical factor for improving model performance (Kumar et al., 2017; Wang et al., 2023; Whang et al., 2023).
+For example, state-of-the-art LLM models, such as ChatGPT (OpenAI, 2022), PaLM (Anil et al., 2023), Gemini (Team et al., 2023) and Mixtral (Jiang et al., 2023) all leverage similar transformer architectures, but they employ distinct data organization and management strategies.
+Therefore, effective training data organization and management has emerged as a decisive factor in scaling model training and enhancing model capabilities (Miao et al., 2024; Fernandez et al., 2023; Prakash et al., 2024; Whang et al., 2023), particularly as models continue to grow in complexity and size.
+The long-context capability is one of the most important features in LLMs (Fu et al., 2024; Xiong et al., 2023; Li et al., 2023), referring to their ability to process and generate extended text while maintaining coherence, retrieving relevant information (Liu et al., 2024), and handling tasks that involve substantial amounts of text efficiently (Bai et al., 2023).
+For instance, GPT-4 (Achiam et al., 2023) can handle context windows of up to 128K tokens, and maintain thematic and logical consistency across long-form text.
+Achieving this long-context capability requires training the model on an extended context window, which usually consist of multiple documents of varying lengths and different domains.
+The primary challenges lie in determining effective data organization and management strategies for long-context model training, which ensure the model achieves long-context capability without compromising its performance in other tasks, such as multitask language understanding (Huang et al., 2019).
+Specifically, the strategy of selecting appropriate documents from vast sources of varying lengths to compose each context window during training remains under-explored, yet it is crucial for maximizing model efficiency and performance.
+Existing methods primarily focus on the combination of document sources, aiming to enhance the model’s general understanding and generation capabilities.
+However, they often overlook the consideration of document length, limiting the model’s ability to effectively handle long-sequence modeling.
+For example, ICLM (Shi et al., 2023) groups related documents within the same context window, enabling LLMs to access a broader range of relevant contexts, while SPLICE (Staniszewski et al., 2023) employs retrieval-based strategies (Robertson et al., 2009; Izacard et al., 2021) to implement structured packing within the context window.
+Unlike previous approaches, we try to explore the data management strategy for long-context model training from a multi-objective combinatorial optimization problem, which involves both document source and length:
+How to design a data organization and management strategy for LLMs that could achieve long-context capability without compromising the models’ performance in other tasks?
+Through extensive experimental analysis, we identified that the challenges associated with data organization and management strategies stem from the following three key aspects:
+C1: Shortage of Long Documents Across Multiple Domains.
+Achieving a balanced domain mixture in training data is essential, as each domain contributes distinct knowledge to the model.
+However, data from various domains exhibit significant variation in length distribution, complicating the creation of effective long-context training datasets.
+While some domains provide sufficiently long documents suitable for long-context training, others, particularly web-sourced data, contain shorter documents, with less than 1% exceeding 4K tokens, as shown in Figure 2.
+To create adequate training contexts from such data, it is often necessary to concatenate multiple documents.
+This process, however, introduces potential challenges, such as increased cross-attention between unrelated documents, leading to undesirable learning signals and higher computational costs.
+Consequently, the variation in document length across domains can hinder the model’s ability to fully leverage long-context capabilities without introducing noise or redundancy.
+C2: Effective Construction for Context Window.
+Developing robust algorithms for constructing context windows is a critical challenge in enhancing the long-context capabilities of language models.
+These algorithms must intelligently group related documents to preserve semantic coherence while maximizing the utility of the available data.
+Poorly designed methods can result in the combination of unrelated or loosely connected documents, which diminishes the quality of training signals and ultimately leads to suboptimal model performance.
+Although existing research has shown the positive impact of structured data on improving language models’ long-context capabilities (Levine et al., 2021; Chan et al., 2022; de Vries, 2023; Devlin et al., 2018), these efforts remain in their infancy.
+Previous approaches (Shi et al., 2023) have frequently resulted in isolated documents, as each document in the training corpus was accessed only once.
+This limitation necessitates the concatenation of unrelated documents with low semantic correlation in subsequent stages, as illustrated inTable 2 and Figure 4(b).
+Moreover, these methods often neglect the essential task of preserving individual document integrity, concentrating on a narrow set of factors rather than adopting a comprehensive framework for context window construction.
+C3: Efficient Organization for Large-Scale Datasets.
+Efficiently organizing large-scale datasets into long-context sequences presents a significant challenge (To et al., 2018; Chen et al., 2024a).
+This task involves not only managing vast amounts of data but also ensuring that the organizational process minimizes computational overhead and labor costs.
+Inefficient data management strategies can lead to increased training times, elevated computational expenses, and difficulties in effectively scaling the model.
+Consequently, optimizing the data organization process is crucial for developing more efficient and scalable large language models (Polyzotis et al., 2017; Pan et al., 2024; Abadi et al., 2009).
+Such strategies must ensure a balance between long-context processing and the practical constraints of resource availability, ultimately leading to improved performance and scalability.
+To address these challenges, we introduce DataSculpt, a novel data management framework designed for long-context training, which formulates the organization of training data as a multi-objective combinatorial optimization (MOCO) problem.
+DataSculpt aims to strike a balance among several key objectives, including maximizing relevance and homogeneity among documents within the same context window (C1, C2), preserving document integrity when truncating text that exceeds the context window’s capacity (C2), and optimizing computational efficiency in handling large-scale datasets (C3).
+Specifically, our approach employs a coarse-to-fine methodology to optimize the organization of training data efficiently and effectively.
+We start by clustering the data based on semantic similarity (coarse), then perform a multi-objective greedy search within each cluster to score and concatenate documents into different context windows (fine).
+To verify the effectiveness of our methods, we conducted continual pretraining with 15 billion tokens on a language model with 7 billion parameters, utilizing varying sequence lengths of 16K, 32K, and 64K.
+This approach enables the model to continuously adapt and refine its knowledge, resulting in enhanced performance over a one-time pretraining method, making it the most widely used.
+Our experimental results demonstrate that our DataSculpt significantly enhances performance across a range of tasks, including general understanding ($+4.88\%$), retrieval augmentation ($+18.09\%$), summarization ($+21.23\%$), reading comprehension ($+21.27\%$), and code completion ($+3.81\%$).
+These findings indicate that DataSculpt effectively synergizes the model’s performance in long-context utilization during the continual pretraining phase, even with a limited number of training tokens.
+Paper Organization.
+In Section 2, we introduce the background concepts relevant to this work, including transformer training and the multi-objective combinatorial optimization problem.
+Then, we present a quantitative analysis of the training datasets in Section 3, emphasizing the challenges faced in long-context LLM training.
+In Section 4, we introduce DataSculpt, which uses a coarse-to-fine methodology for efficiently organizing training data, enabling the model to achieve long-context capability without compromising performance on other tasks.
+Finally, we validate the effectiveness of DataSculpt through comprehensive experiments in Section 5.
+2. Preliminary
+2.1. Transformer Architecture
+Transformers are designed to address sequence modeling and transduction tasks, such as language modeling and machine translation (Vaswani et al., 2017).
+At the heart of their architecture lie self-attention networks and point-wise feed-forward networks, which are integral to each layer.
+These models perform dense algebraic operations, including matrix multiplications, which incurs significant computational and memory demands.
+Self-Attention Network.
+Self-attention stands for a critical component of the Transformer architecture, allowing the model to weigh the importance of different words in a sequence, irrespective of their positions (Zhang et al., 2019; Miao et al., 2022; Vaswani et al., 2017).
+For each word in a sequence, self-attention computes a score that represents how much focus to place on other parts of the input sequence when encoding that specific word.
+This mechanism involves three sets of weights: query, key, and value, which are learned during training.
+Mathematically, self-attention can be described as a mapping of a query and a set of key-value pairs to an output, where the output is a weighted sum of the values.
+Each weight assigned to a value is computed by a compatibility function of the query with the corresponding key.
+This allows the Transformer to capture complex dependencies and relationships within the data, making it highly effective for tasks that require an understanding of the entire sequence, such as text translation or summarization.
+Feed-Forward Network.
+Each transformer layer includes a position-wise feed-forward network (FFN), which is applied to each position separately and identically.
+This consists of two linear transformations with a ReLU activation in between.
+The purpose of the FFN is to transform the representation independently at each position with the same, learned linear transformation.
+This means that while the self-attention layers share information across positions in the input, the feed-forward network processes each position independently.
+This component is crucial as it introduces non-linearity into the model, enabling it to learn more complex functions.
+Moreover, it ensures that the Transformer can adapt to different data types and tasks, providing additional flexibility and power to the model (Vaswani et al., 2017).
+2.2. LLMs Training
+Large Language Models (LLMs) denote Transformer-based architectures equipped with hundreds of billions (or more) of parameters (Zhao et al., 2023), initially introduced to address sequence modeling and transduction challenges.
+Below, we have outlined several key concepts about LLMs training.
+Pretraining.
+The pretraining phase of LLMs is a crucial period for developing their general knowledge capabilities.
+During this phase, LLMs undergo unsupervised training with large-scale raw text to cultivate their language modeling skills.
+Given a sequence $\mathrm{\textbf{x}}=\{x_{1},x_{2},...,x_{n}\}$, the primary training task for LLMs at this stage is to predict subsequent tokens $x_{i}$ based on preceding tokens $\mathrm{\textbf{x}}_{<i}$ as context, outlined in Fig.
+1.
+The objective function aims to maximize the likelihood of Equation 1.
+(1) $\mathcal{L}_{LLM}^{PT}(\mathrm{\textbf{x}})=\sum_{i=1}^{n}\log{P(x_{i}|\mathrm{\textbf{x}}_{<i})}$
+Continual Pretraining.
+Continual pretraining is a phase that extends the initial pretraining, using the same objective as shown in Equation 1.
+The key distinction is that continual pretraining builds on an already trained model rather than starting from scratch.
+At this stage, the LLM has basic language modeling capabilities; the aim now is to enhance its specialized functions, such as long context modeling, leveraging the foundational skills acquired during the initial pretraining phase.
+This phase is increasingly seen as a practical approach to improving model performance without the exorbitant computational costs associated with starting anew, as argued in recent studies (Ibrahim et al., 2024; Gupta et al., 2023; Parmar et al., 2024).
+Context Window.
+The context window is the span of text that a model actively considers when predicting the next word or deciphering the text’s structure.
+This window defines the range of contextual neighbors used in calculating vector representations (Lison and Kutuzov, 2017).
+During both the pretraining and continual pretraining phases, raw texts from various documents are concatenated to form a continuous sequence that matches the length of the context window.
+Any text exceeding this length is truncated.
+The size of the context window is pivotal, as it directly impacts the model’s capacity for language comprehension and generation.
+Figure 1. An illustration of the data processing workflow, training task, context window and self-attention mechanism.
 Figure 2.
-Distribution of token counts for individual documents across different domains.
-Domains are arranged in descending order from top to bottom based on the proportion of documents with token numbers exceeding 64K.
-Documents within the domains of books and academic papers predominantly feature medium to long lengths (L $\geq$ 16K), in contrast to other fields, especially web-sourced data, where document lengths are almost exclusively below 4K.
-C1.
-Insufficient Domain-Balanced Long Training Corpora.
-The challenge lies in effectively incorporating a wide range of domains into extended-context LLM training.
-Achieving a balanced domain mixture is imperative, since each domain encompasses distinct knowledge.
-However, the distribution of text lengths varies substantially across various domains, as depicted in Figure 1.
-While some domains offer data lengths suitable for long-context training, other domains, particularly web-sourced data where less than 1% of documents exceed 4K tokens, necessitate the concatenation of multiple documents to achieve adequate training contexts.
-This may lead to increased cross-attention among different documents within sequences, which is an undesirable learning signal.
-C2.
-Absence of Criteria for Long Context Construction.
-While existing work has demonstrated the positive impact of structured data on amplifying language models’ long-context capabilities (Levine et al., 2021; Chan et al., 2022; de Vries, 2023a; Devlin et al., 2018), these efforts remain preliminary.
-Research dedicated to the organization and quality assessment of long-sequence training data for LLMs tends to focus on a narrow set of factors, lacking a systematic framework.
-Consequently, there is a discernible need for a detailed and cohesive framework that systematically organizes training data for applications requiring extensive context.
-C3.
-Efficiency in Massive Data Management.
-Training LLMs with extended contexts from scratch is notably resource-intensive, requiring considerable time and labor investments (Chen et al., 2024a), which has redirected emphasis towards post-training optimizations as a more feasible strategy.
-Furthermore, organizing extensive pretraining datasets into long-context sequences poses substantial logistical hurdles (To et al., 2018), further compounding the complexity of developing LLMs efficiently.
-Consequently, devising and implementing strategies for efficient data scheduling and management is essential to enhance the capabilities of LLMs in a sustainable and cost-effective manner (Polyzotis et al., 2017), balancing the need for extensive context processing with the practical limitations of resource availability.
-To address these issues, we introduce DataSculpt, a novel framework designed to reconfigure the organization of extended-context training data into a multi-objective combinatorial optimization problem, as illustrated in Figure 2.
-DataSculpt meticulously balances several critical objectives throughout the data organization process, including relevance, homogeneity, integrity, as well as computational efficiency, adopting a coarse-to-fine methodology to optimize the architecture of training data.
-We conducted post-training with 15 billion tokens on a 7B model across varying sequence lengths of 16K, 32K, and 64K, and experimental results demonstrate that DataSculpt improves performance in an array of tasks, most significantly in general understanding (+4.88%), retrieval augmentation (+18.09%), summarization (+21.23%), reading comprehension (+21.27%) and code completion (+3.81%).
-The finding reveals that DataSculpt can markedly synergize the model’s performance in long-context utilization during the post-train phase, even with a constrained number of training tokens.
-Our contributions can be summarized as follows:
-  • We highlight the challenge of compiling an appropriate long-context training corpus with a balanced domain mixture, crucial for enhancing the capabilities of LLMs in processing extensive text sequences.
-  • To create suitable long context training data, we propose an efficient framework DataSculpt, which employs a multi-objective partition allocation strategy for the construction of extended context sequences.
-  • Through quantitative analysis and t-SNE visualization of clustering outcomes at the intermediary stage of DataSculpt, we validate DataSculpt’s robustness for multi-objective data concatenation within extended sequences.
-  • We engaged in continual pretraining using DataSculpt with various context lengths on a 7B model, and subsequently assessed its performance across 17 distinct tasks, illuminating the profound impact of our proposed method.
-2. Related Work
-2.1. Training Long-Context Language Models
-Training Large Language Models (LLMs) with expanded context windows facilitates the execution of tasks that surpass the capabilities defined by the existing paradigm, such as multi-document question answering (Caciularu et al., 2023), long document summarization (Bai et al., 2023), and enhanced retrieval functionalities (Liu et al., 2024).
-Nevertheless, the endeavor of training LLMs with extended contexts is notably time-consuming and labor-intensive.
-Furthermore, documents with long sequences are conspicuously scarce in the training corpus (de Vries, 2023a).
-For instance, less than 5% of documents within CommonCrawl exceed 2k tokens in length (Cafarella and Halevy, 2011), which hinders LLMs’ exposure to complex narratives.
-Despite these challenges, significant strides have been made toward enhancing LLMs’ capabilities for long context processing.
-Architecture-Based Strategies.
-Numerous studies explored the variants of Transformer architecture by modifying its attention mechanism, which includes introducing recurrence (Dai et al., 2019), transforming attention into less computationally demanding forms (Beltagy et al., 2020; Zaheer et al., 2020), and employing low-rank approximations (Wang et al., 2020; Peng et al., 2021).
-Furthermore, LONGLLAMAs (Tworkowski et al., 2024) presented the Focused Transformer (FOT) featuring a novel training objective that enhances the (key, value) structure, which facilitates the expansion of context length and enables efficient scaling through k-nearest neighbors lookup.
-These approaches achieve more efficient exact attention through the development of an IO-optimized CUDA kernel specifically designed for this purpose.
-Data-Centric Approaches.
-The effectiveness of long-context language models relies equally upon the innovative employment of data-centric strategies throughout their training processes.
-The construction of long-context data involves more than a mere concatenation of short documents, which can lead to cross-document attention within a sequence.
-In these circumstances, formulating high-quality or domain-specific long sequences holds significant research value (de Vries, 2023b).
-ICLM (Shi et al., 2023) and SPLICE (Staniszewski et al., 2023) combine similar training instances through a sophisticated retrieval process, which emphasizes the critical role of high-quality long-context training data in enhancing the context expansion capabilities of LLMs without necessitating extensive data resources.
-Additionally, Krell et al.
-(Krell et al., 2021) deliberate on the challenges associated with the conventional concat-and-chunk approach, proposing an approximate bin-packing algorithm as a potential solution.
-As an improvement, dataset decomposition (Pouransari et al., 2024) and Best-fit Packing (Ding et al., 2024) both propose the sequence length aware optimization technique to minimize truncations, thereby facilitating more efficient pretraining on long sequences and achieving scalability relative to the size of dataset.
-Moreover, ProLong (Chen et al., 2024a) introduces an evaluation framework for assessing the quality of long-context data, further enriching the field’s methodological toolkit.
-Nonetheless, these methods have not performed a systematic framework tailored for data construction in long context training.
-Our work conceptualizes data management for extended-context training as a multi-objective combinatorial optimization problem, and provide a comprehensive framework for long sequence data organization.
-2.2. Multi-Objective Combinatorial Optimization (MOCO)
-Multi-objective combinatorial optimization (MOCO) problems are characterized by the need to optimize multiple objectives simultaneously, which are often in conflict with one another (Ulungu and Teghem, 1994).
+An illustration on the shortage of long documents across multiple domains.
+Left: The length distribution of individual documents across various domains, arranged in descending order by the proportion of documents exceeding 64K tokens.
+Documents, such as books and academic papers predominantly conist of medium to long lengths (L $\geq$ 16K), whereas other domains, particularly web-sourced data, contain documents almost exclusively shorter than 4K tokens.
+Right: The proportion of each domain in the overall training dataset.
+Table 1. Token statistics across domains. We analyzed the maximum, minimum, and average token counts for different domains.
+Domain Book (Chinese) Book (English) Paper (English) Code Paper (Chinese) Web (English) Web (Chinese) Others
+Max Len 24,203,838 8,241,838 10,721,906 1,000,000 1,471,297 1,267,049 80,688 21,839
+Min Len 667 322 127 173 696 289 132 23
+Avg Len 176,156 216,969 20,555 2,578 6,216 886 358 224
+2.3. Multi-Objective Combinatorial Optimization
+The multi-objective combinatorial optimization (MOCO) problem involves optimizing several objectives simultaneously, which typically conflict with each other (Ulungu and Teghem, 1994) and can be formulated as:
+$\max_{x\in\chi}\mathbf{f}\left(x\right)$ $:=\left(f_{1}\left(x\right),f_{2}\left(x\right),...,f_{m}\left(x\right)\right)$
+(2) s.t. $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k},k=1,...,n$
+where $x$ denotes a solution in the decision space $\chi\subset\mathbb{R}^{n}$, $\mathbf{f}:\chi\to\mathbb{R}^{m}$ represents an $m$-dimensional vector-valued objective function, and $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k}\,\text{where}\,k=1,...,n$ stands for a series of soft constraints.
+Importantly, the evaluation process is computationally expensive for each individual objective $f_{i}\left(x\right),i=1,...,m$.
 The complexity of MOCO challenges lies in the fact that improving the outcome of one objective often leads to the deterioration of another, making it necessary to find a balance or trade-off between them.
-MOCO-Augmented LLMs.
-Existing works have integrated MOCO problems with relevant concerns in the LLMs field.
-Wang et al.
-(Wang et al., 2024) employed the MOCO approach within the Reinforcement Learning from Human Feedback (RLHF) framework for LLMs, where the reward models aggregate multiple reward objectives with coefficients into a scalar score.
-Akiba et al.
-(Akiba et al., 2024) viewed model merging as a MOCO challenge, and introduced an innovative approach that leverages evolutionary algorithms to facilitate the automated development of robust foundation models.
-Furthermore, integration of MOCO with data-centric LLMs have been explored for training data selection (Du et al., 2023; Xu et al., 2023; Chen et al., 2024a), which applies heuristic techniques to assign weights to different data evaluation metrics and facilitates a phased resolution approach.
-Considering the extensive scale of training dataset, data construction for long context is treated as a MOCO challenge, for which the coarse-to-fine algorithm is identified as a suitable resolution strategy.
-3. Method
-We propose an innovative approach to construct concatenated sequences with multiple attributes for the enhancement of LLMs long-context utilization.
-3.1. Multi-Objective Partition Allocation
+3. Observations and Challenges
+3.1. Observations on Training Datasets
+The training data is sourced from a variety of domains, such as web content, books, academic papers, and code repositories, with the composition and respective weight proportions shown on the right side of Figure 2.
+We conducted a statistical analysis of token count characteristics across various domains, as shown in Table 1.
+Additionally, we categorized tokenized documents into different length intervals based on the token counts of each data source.
+The proportional distributions across domains are visualized as a normalized stacked horizontal bar chart on the left size of Figure 2, offering a hierarchical view of the domains according to the prevalence of longer texts.
+From Table 1 and Figure 2, we have observed that data from the book and academic paper domains predominantly consist of long documents, with over half exceeding 64K tokens in length.
+Academic papers, however, tend to fall within the medium-length range, with documents typically spanning 4K to 32K tokens.
+In contrast, web-sourced data is primarily composed of short documents, with the vast majority under 4K tokens.
+As a result, web-sourced data is not naturally suited for long-context training due to its shorter document lengths, which often requires the concatenation of multiple documents to form sequences of adequate length for long-context training.
 Figure 3.
@@
 During the semantic clustering phase, we implement a variant of the ISODATA algorithm, which is augmented with the FAISS vector searching library, to aggregate documents based on semantic similarity.
-Following this, a greedy semantic-driven largest-fit algorithm is utilized to arrange these clusters into sequences that are multi-objective optimally configured for extended-context training.
-We consider the problem as a multi-objective partition allocation problem with soft constraints in a massive data scale, which is essentially an instance of the multi-objective combinatorial optimization (MOCO) problem (Lin et al., 2022).
-This can be formulated as
-$\max_{x\in\chi}\mathbf{f}\left(x\right)$ $:=\left(f_{1}\left(x\right),f_{2}\left(x\right),...,f_{m}\left(x\right)\right)$
-(1) s.t. $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k},k=1,...,n$
-where $x$ denotes a solution in the decision space $\chi\subset\mathbb{R}^{n}$, $\mathbf{f}:\chi\to\mathbb{R}^{m}$ represents an $m$-dimensional vector-valued objective function, and $g_{k}\left(x\right)\leq b_{k}+\epsilon_{k}\,\text{where}\,k=1,...,n$ stands for a series of soft constraints.
-Importantly, the evaluation process is computationally expensive for each individual objective $f_{i}\left(x\right),i=1,...,m$.
+Following this, a greedy semantic-driven largest-fit algorithm is utilized to arrange these clusters into sequences that are multi-objective optimally configured for long-context training.
+3.2. Challenges
+Given the token count statistics across documents, one of the primary challenges lies in the variation of sequence lengths among different data sources.
+In efforts to preserve domain diversity, shorter documents in the training dataset tend to increase the incidence of cross-document attention, where the model focuses on multiple, unrelated documents within the same context window.
+This can lead to model hallucinations and increased computational consumption.
+The problem becomes more prominent as the model’s context window expands during training.
+Model Hallucinations.
+Model hallucinations occur when models generate incorrect or irrelevant content that is not grounded in the input data.
+While well-trained models generally reduce cross-document attention, this behavior is not always strictly controlled, which may potentially lead to spurious outputs as well as unintended correlations (Pouransari et al., 2024; Ding et al., 2024).
+Computational Consumption.
+Moreover, cross-document attention expends unnecessary computational resources by focusing on unrelated tokens that contribute little to the model’s learning process (Krell et al., 2021; Variš and Bojar, 2021).
+The unnecessary processing of unrelated tokens not only diminishes the model’s performance but also amplifies the computational demands, which is especially crucial given the attention mechanism’s quadratic complexity (Vaswani et al., 2017).
+3.3. Problem Formulation
+Based on the above observations and analysis, we frame the scenario of data organization and management for long-context training as a multi-objective partition allocation problem with soft constraints at a massive data scale.
+This is fundamentally an instance of the multi-objective combinatorial optimization (MOCO) problem (Lin et al., 2022) described in Section 2.3.
 Setup.
-Suppose that we have a training corpus $\mathcal{D}$ and training sequences $\mathcal{W}$ pending sampling, the context window for each is limited in length $L$.
-The multi-objective partition allocation problem encompasses the following optimization objectives, the ultimate goal of which is to efficiently amplify the homogeneity of concatenated documents within the identical attention sequence:
+Given a training dataset $\mathcal{D}$ and training sequences $\mathcal{W}$ awaiting sampling, with each context window limited to a length of $L$, the multi-objective partition allocation problem focuses on optimizing several key objectives.
+The overarching goal is to efficiently enhance the homogeneity of concatenated documents within the same attention sequence:
   • G1: Relevance.
-One fundamental goal is to maximize the holistic semantic coherence among documents contained in the same context window, with the observation that indicates a reduction in the perplexity of language models with increasing relevance (Shi et al., 2023).
+The primary goal is to maximize semantic coherence among documents within the same context window.
+Increasing relevance has been shown to reduce language model perplexity, improving overall model performance (Shi et al., 2023).
   • G2: Homogeneity.
-Another objective to be pursued is diminishing cross-document attention during the model’s training phase, which can be accomplished by mitigating the quantity of documents utilized for concatenation within each sequence.
-  • G3: Integrity.
-Eliminating unnecessary truncation when concatenating documents for each training sequence is essential for preserving the integrity of individual documents (Ding et al., 2024), ensuring that the contextual richness and informational completeness are maintained to the greatest extent possible.
-  • G4: Compute Efficiency.
-In the process of managing vast training datasets for LLMs, efficiency emerges as a critical consideration.
-When aligning long-context training data with multi-objective goals, a balance among accuracy, processing speed, and space consumption is vital for enhancing LLM performance while optimizing resource efficiency.
-Soft constraints mandate that the sum of document length within each individual context window can not surpass the respective capacity, with the aim of maximizing document integrity and minimizing truncated token proportion within the same document (G3).
-Thus, Equation 1 can be written as
+Reducing cross-document attention during the model’s training phase is crucial.
+This can be achieved by minimizing the number of concatenated documents within each context window, ensuring the model focuses on fewer, more cohesive sequences.
+  • G3: Integrity.Preserving the integrity of individual documents by minimizing unnecessary truncation during concatenation is essential.
+This ensures that the contextual richness and completeness of each document are retained as much as possible (Ding et al., 2024).
+  • G4: Efficiency.Efficiency is a key concern when managing large-scale training datasets for LLMs.
+Achieving a balance between accuracy, processing speed, and space consumption is critical for optimizing long-context training performance while minimizing resource usage.
+Soft constraints ensure that the total length of documents within each context window does not exceed its capacity, aiming to maximize document integrity and minimize the proportion of truncated tokens within the same document (G3).
+Consequently, Equation 2 can be written as:
 (3) $\max_{x\in\chi}\mathbf{f}\left(x\right):=\alpha f_{1}\left(x\right)+\beta f_{2}\left(x\right)+\lambda p\left(x\right)$
-where soft constraints are transformed into penalty terms, while $\alpha,\beta,\lambda$ are non-negative hyper-parameters that control the impact of objective and penalty functions.
-The respective objective functions are
+where soft constraints are incorporated as penalty terms, while $\alpha$, $\beta$, $\lambda$ are non-negative hyper-parameters that control the weight of each objective and penalty function.
+The respective objective functions can be formulated as:
 (4) $f_{1}\left(x\right)$ $=\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}\sum_{\begin{subarray}{c}d_{j}\in\mathcal{D}\\ j\neq i\end{subarray}}x_{d_{i}w}x_{d_{j}w}\cos\left(E(d_{i}),E(d_{j})\right)$
@@
 $L$ is used to divide the objective function $f_{2}\left(x\right)$ to transform the minimization problem into a maximization problem for resolution.
-Moreover, the penalty term $p(x)$ can be written as
-(4) $p(x)=\begin{cases}0,\text{if }\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}\leq L&\\ \frac{L}{\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}-1,\text{otherwise}&\end{cases}$
-which serves as a penalty for the truncation of concatenated documents that exceed context window capacity, thereby causing document incompleteness.
-Solving MOCO problems exactly is NP hard, but coarse-to-fine algorithms are recognized for offering an efficient approximate solution (G4).
-However, the challenge lies in effectively reconciling multiple objects within the constraints of computational resources and the vast scale of the training corpus $\mathcal{D}$ (Zhao et al., 2005).
-We denote $|\mathcal{D}|$ as the number of documents contained in $\mathcal{D}$.
-The quantity of documents $k$ chosen for concatenation within a specific context is notably smaller than the total document count in $\mathcal{D}\,(k<<|\mathcal{D}|)$.
-Consequently, despite conducting score-based sorting on the remaining $|\mathcal{D}|-k$ documents during each iteration, they remain unselected, thereby consuming a significant expenditure of time and computational resources.
-Therefore, we initially perform semantic clustering on the entire training corpus $\mathcal{D}$ at a coarser granularity, followed by the implementation of a greedy algorithm to optimize the trade-off between efficiency and precision during the document organization process.
-3.2. Phase 0: Data Preprocessing
-Given the training context window length $L$ of the model, we tokenize raw documents in the training corpus $\mathcal{D}$ and divide them into segments, each comprising no more than $L$ tokens, which is dictated by the context length constraints.
-Subsequently, we utilize the embedding model to embed each segmented document $d_{i}$ to get $E\left(d_{i}\right)$.
+Furthermore, the penalty term $p(x)$ is defined as:
+(5) $p(x)=\begin{cases}1,\text{if }\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}l(d_{i})\leq L&\\ \frac{L}{\sum_{w\in\mathcal{W}}\sum_{d_{i}\in\mathcal{D}}x_{d_{i}w}}-1,\text{otherwise}&\end{cases}$
+where $l(d_{i})$ represents the length of document $d_{i}$.
+The penalty term ensures that the total length of concatenated documents remains within the capacity $L$ of the context window.
+If the total sequence length exceeds $L$, the penalty increases to discourage overfitting the context window.
+Solving MOCO problems exactly is NP-hard, but coarse-to-fine algorithms are known for providing efficient approximate solutions (G4).
+The primary challenge, however, lies in effectively balancing multiple objectives within the constraints of limited computational resources and the vast scale of the training corpus $\mathcal{D}$ (Zhao et al., 2005).
+Let $|\mathcal{D}|$ denote the total number of documents in $\mathcal{D}$.
+The number of documents $k$ selected for concatenation within a given context window is significantly smaller than the total document count, such that $k<<|\mathcal{D}|$.
+As a result, even though score-based sorting is performed on the remaining $|\mathcal{D}|-k$ documents during each iteration, they remain unselected, consuming substantial time and computational resources.
+To address this inefficiency, we first apply semantic clustering to the entire training corpus $\mathcal{D}$ at a coarser granularity.
+This is followed by a greedy algorithm that optimizes the trade-off between computational efficiency and precision during the document organization process.
+4. DataSculpt
+In this work, we introduce DataSculpt, a novel data management framework designed for long-context training.
+This framework aims to effectively integrate data from multiple domains and optimize the context window during training.
+Figure 3 presents the workflow of DataSculpt.
+In the data preprocessing phase, we first tokenize the raw documents from the training corpus, $\mathcal{D}$, according to the context window length, $L$, of the model.
+This process involves dividing the documents into segments, each containing no more than $L$ tokens, to comply with the model’s context length constraints.
+Following this, each document segment $d_{i}$ is processed through the embedding model to generate its embedding $E(d_{i})$.
 4.1. Phase 1: Semantic Clustering
+Input: Training corpus $\mathcal{D}$, Cluster number $N_{c}$ gained from Equation 7, Hyperparameters: distance threshold $\delta$, alteration threshold $\epsilon$, maximum iterations $T$;
+Parameter: Unclustered data $\mathcal{U}$, Alteration distance $dst$;
+Ouptut: Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$;
+Define: $\mu(c)$ is center vector of cluster $c$;
+1 Sample $N_{c}$ centroids in $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ randomly from $\mathcal{D}$, $\mu(c_{j})\leftarrow E(d_{j})\,\text{for all }j=1,...,N_{c}$ ;
+[…]
+3 for $t=1,2,\ldots,T$ do
+4       for each documnet $d_{i}$ in $\mathcal{D}$ do
+5             if $\max_{c_{j}\in\mathcal{C}}cos\left(E\left(d_{i}\right),\mu\left(c_{j}\right)\right)>\delta$ or $t==T$ then
+6                   Add $d_{i}$ to $c_{j^{*}}=\arg\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))$;
+7
+8             $\mathcal{U}.append(d_{i})$;
+9
+10      for each $d_{u}$ in $\mathcal{U}$ do
+11             Add new cluster $c_{u}$, $\mu(c_{u})\leftarrow E(d_{u})$, $N_{c}\leftarrow N_{c}+1$;
+12
+13      for each cluster $c_{j}$ do
+14             $dst\leftarrow dst+||\mu(c_{j})-\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})||$;
+15             Recalculate centroid $\mu(c_{j})\leftarrow\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})$;
+16
+17      for each pair of clusters $c_{i}$ and $c_{j}$ do
+18             if $cos(\mu(c_{i}),\mu(c_{j}))>\delta$ then
+19                   Merge cluster $c_{i}$ and cluster $c_{j}$, $N_{c}\leftarrow N_{c}-1$;
+20
+21      if $dst<\epsilon$ then
+22             Break;
+23
+24      $dst\leftarrow 0$;
+25
+26return Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ where $\cup_{j=1}^{N_{c}}c_{j}=\mathcal{D}$ ;
 Algorithm 1 Variant of ISODATA
-Input: Training corpus $\mathcal{D}$, Cluster number $N_{c}$ gained from Equation 6, Hyperparameters: distance threshold $\delta$, alteration threshold $\epsilon$, maximum iterations $T$
-Parameter: Unclustered data $\mathcal{U}$, Alteration distance $dst$
-Output: Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$
-Define $\mu(c)$: center vector of cluster $c$
-1:  Sample $N_{c}$ centroids in $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ randomly from $\mathcal{D}$, $\mu(c_{j})\leftarrow E(d_{j})\,\text{for all }j=1,...,N_{c}$
-[…]
-3:  for $t=1,2,\ldots,T$ do
-4:     for each documnet $d_{i}$ in $\mathcal{D}$ do
-5:        if $\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))>\delta$ or $t==T$ then
-6:           Add $d_{i}$ to $c_{j^{*}}=\arg\max_{c_{j}\in\mathcal{C}}cos(E(d_{i}),\mu(c_{j}))$
-7:        else
-8:           $\mathcal{U}.append(d_{i})$
-9:        end if
-10:     end for
-11:     for each $d_{u}$ in $\mathcal{U}$ do
-12:        Add new cluster $c_{u}$, $\mu(c_{u})\leftarrow E(d_{u})$, $N_{c}\leftarrow N_{c}+1$
-13:     end for
-14:     for each cluster $c_{j}$ do
-15:        $dst\leftarrow dst+||\mu(c_{j})-\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})||$
-16:        Recalculate centroid $\mu(c_{j})\leftarrow\frac{1}{|c_{j}|}\sum_{d_{i}\in c_{j}}E(d_{i})$
-17:     end for
-18:     for each pair of clusters $c_{i}$ and $c_{j}$ do
-19:        if $cos(\mu(c_{i}),\mu(c_{j}))>\delta$ then
-20:           Merge cluster $c_{i}$ and cluster $c_{j}$, $N_{c}\leftarrow N_{c}-1$
-21:        end if
-22:     end for
-23:     if $dst<\epsilon$ then
-24:        Break
-25:     end if
-26:     $dst\leftarrow 0$
-27:  end for
-28:  return  Clusters $\mathcal{C}=\left\{c_{j}\right\}_{j=1}^{N_{c}}$ where $\cup_{j=1}^{N_{c}}c_{j}=\mathcal{D}$
-In the initial phase, the documents undergo a process of semantic clustering.
+In the initial phase, the documents undergo a process of semantic clustering at a relatively coarse granularity.
 To estimate the initial cluster number, we derive inspiration from the construction of a distance matrix by quantifying semantic distances between documents within the training corpus $\mathcal{D}$.
@@
 Furthermore, to enhance the efficiency of vector similarity assessments, we have integrated the FAISS index (Johnson et al., 2019), a sophisticated method for conducting high-speed pairwise similarity searches.
-We first randomly select $N_{c}$ data samples from the entire training corpus $\mathcal{D}$ to serve as the initial clustering centers, which is based on the calculations from Equation 6.
+As shown in Algorithm 1, we first randomly select $N_{c}$ data samples from the entire training corpus $\mathcal{D}$ to serve as the initial clustering centers (Line 1), which is based on the calculations from Equation 7.
 These sampled data points are then indexed using the FAISS vector searching library.
-Throughout the clustering process, for each document $d_{i}$, we employ the HNSW method (Malkov and Yashunin, 2018) from FAISS to retrieve its nearest cluster center $c_{j}$.
-If the semantic similarity between the current center and $d_{i}$ exceeds a predefined threshold $\delta$, then $d_{i}$ is assigned to cluster $c_{j}$.
-Conversely, if the similarity does not surpass $\delta$, $d_{i}$ is designated as a new cluster center.
-Moreover, as the iterative clustering progresses, adjustments to the cluster centers are made by calculating the mean of the document embedding vectors within each cluster, thereby establishing new cluster centroids.
-Additionally, if the similarity between any two clusters exceeds the threshold $\delta$, those clusters are merged.
+Throughout the clustering process with a maximum iteration count of $T$ (Line 3-18), for each document $d_{i}$, we employ the HNSW method (Malkov and Yashunin, 2018) from FAISS to retrieve its nearest cluster center $c_{j}$ (Line 4-7).
+If the semantic similarity between the current center and $d_{i}$ exceeds a predefined threshold $\delta$, then $d_{i}$ is assigned to cluster $c_{j}$ (Line 5-6).
+Conversely, if the similarity does not surpass $\delta$, $d_{i}$ is designated as a new cluster center (Line 7).
+Moreover, as the iterative clustering progresses, adjustments to the cluster centers are made by calculating the mean of the document embedding vectors within each cluster, thereby establishing new cluster centroids (Line 10-12).
+Additionally, if the similarity between any two clusters exceeds the threshold $\delta$, those clusters are merged (Line 13-15).
 This approach not only refines the clustering process by dynamically adjusting cluster centers based on semantic coherence but also ensures that clusters remain distinct and meaningful by merging overly similar clusters, thereby enhancing the overall structure and interpretability of the clustering outcome.
@@
 The purpose of this algorithm, as depicted in Algorithm 2, is to systematically arrange documents into various context sequences for concatenation.
-Algorithm 2 Greedy Semantic-Driven Largest-Fit Allocation
-Input: Cluster $c_{j}\in\mathcal{C}$, Context sequence capacity $L$, Context sequence quantity $N_{w}$, Hyperparameters: impact of objective and penalty function $\alpha,\,\beta,\,\lambda$
-Parameter: Available context sequence $\mathcal{A}$, Semantic relevance $f_{1}$, Residual capacity proportion $f_{2}$, Truncation penalty $p$, Concatenation score $F$
-Output: Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$
-Define $r(s)$: remaining capacity of context sequence $s$
-Define $l(d)$: token quantity of document $d$
-Define $\mu(s)$: center vector of context sequence $s$
-1:  $\left\{r(s_{l})\leftarrow L\right\}_{l=1}^{N_{w}}$: Initialize empty context sequences
-2:  $S_{c_{j}}$: Sort $d_{i}\in c_{j}$ by token quantity in descending order
-3:  for each documnet $d_{i}$ in $S_{c_{j}}$ do
-4:     $\mathcal{A}=\left\{s_{l}|r(s_{l})\geq 0\right\}_{l=1}^{N_{w}}$
-5:     if $\mathcal{A}=\emptyset$ then
-6:        Break
-7:     end if
-8:     for each context sequence $s_{a}$ in $\mathcal{A}$ do
-9:        $f_{1}(d_{i},s_{a})=cos(E(d_{i}),\mu(s_{a}))$
-10:        $f_{2}(d_{i},s_{a})=r(s_{a})/L$
-11:        $p(d_{i},s_{a})=\begin{cases}0,\text{if }l(d_{i})\leq r(s_{a})&\\ L/(L+l(d_{i})-r(s_{a})),\text{otherwise}&\end{cases}$
-12:        $F(d_{i},s_{a})=\alpha f_{1}(d_{i},s_{a})+\beta f_{2}(d_{i},s_{a})+\lambda p(d_{i},s_{a})$
-13:        Add $d_{i}$ to $s_{a^{*}}$ where $s_{a^{*}}=\arg\max_{s_{a}\in\mathcal{S}}F(d_{i},s_{a})$
-14:     end for
-15:     $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$
-16:     $\mu(s_{a^{*}})\leftarrow((|s_{a^{*}}|-1)\mu(s_{a^{*}})+E(d_{i}))/|s_{a^{*}}|$
-17:  end for
-18:  return  Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$
-As delineated in Section 3.1, our objective encompasses augmenting the relevance (G1), homogeneity (G2), and integrity (G3) of the textual content encapsulated within each context window.
+Input: Cluster $c_{j}\in\mathcal{C}$, Context sequence capacity $L$, Context sequence quantity $N_{w}$, Hyperparameters: impact of objective and penalty function $\alpha,\,\beta,\,\lambda$;
+Parameter: Available context sequence $\mathcal{A}$, Semantic relevance $f_{1}$, Residual capacity proportion $f_{2}$, Truncation penalty $p$, Concatenation score $F$;
+Ouptut: Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$;
+Define: $r(s)$ is remaining capacity of context sequence $s$;
+Define: $l(d)$ is token quantity of document $d$;
+Define: $\mu(s)$ is center vector of context sequence $s$;
+1 $\left\{r(s_{l})\leftarrow L\right\}_{l=1}^{N_{w}}$: Initialize empty context sequences;
+2 $S_{c_{j}}$: Sort $d_{i}\in c_{j}$ by token quantity in descending order;
+3 for each documnet $d_{i}$ in $S_{c_{j}}$ do
+4       $\mathcal{A}=\left\{s_{l}|r(s_{l})\geq 0\right\}_{l=1}^{N_{w}}$;
+5       if $\mathcal{A}=\emptyset$ then
+6             Break;
+7
+8      for each context sequence $s_{a}$ in $\mathcal{A}$ do
+9             $f_{1}(d_{i},s_{a})=cos(E(d_{i}),\mu(s_{a}))$;
+10             $f_{2}(d_{i},s_{a})=r(s_{a})/L$;
+11             $p(d_{i},s_{a})=\begin{cases}1,\text{if }l(d_{i})\leq r(s_{a})&\\ L/(L+l(d_{i})-r(s_{a})),\text{otherwise}&\end{cases}$;
+12             $F(d_{i},s_{a})=\alpha f_{1}(d_{i},s_{a})+\beta f_{2}(d_{i},s_{a})+\lambda p(d_{i},s_{a})$;
+13             Add $d_{i}$ to $s_{a^{*}}$ where $s_{a^{*}}=\arg\max_{s_{a}\in\mathcal{S}}F(d_{i},s_{a})$;
+14
+15      $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$;
+16       $\mu(s_{a^{*}})\leftarrow((|s_{a^{*}}|-1)\mu(s_{a^{*}})+E(d_{i}))/|s_{a^{*}}|$;
+17
+18return Constructed sequences $\mathcal{S}=\left\{s_{l}\right\}_{l=1}^{N_{w}}$;
+Algorithm 2 Semantic-Driven Largest-Fit Allocation
+As delineated in Section 3.3, our objective encompasses augmenting the relevance (G1) and homogeneity (G2), while maintaining integrity (G3) of the textual content encapsulated within each context window.
 This enhancement is realized through a strategic reduction in the quantity of documents concatenated into a single context sequence, amplifying the semantic congruence among these documents, and eliminating unnecessary truncations.
-Initially, documents within each previously formed cluster are organized in descending order by their token count.
-Following this organization, for each document $d_{i}$ pending allocation, the process seeks to integrate it into an existing sequence $s_{a}$ with content length that does not exceed the maximum context window capacity $L$.
-The sequence $s_{a}$ is selected based on achieving the highest composite score $F(d_{i},s_{a})$, which reflects a weighted consideration of semantic similarity (G1), available residual capacity (G2), and the minimization of truncation penalties (G3).
-Through this approach, we aim to ensure that the allocation of documents to sequences is not only computing efficient but also optimally maintains semantic consistency, maximizes capacity utilization, and preserves the integrity of individual documents within each training sequence.
+Initially, documents within each previously formed cluster are organized in descending order by their token count (Line 2).
+Following this organization, for each document $d_{i}$ pending allocation, the process seeks to integrate it into an existing sequence $s_{a}$ with content length that does not exceed the maximum context window capacity $L$ (Line 3-14).
+The sequence $s_{a}$ is selected based on achieving the highest composite score $F(d_{i},s_{a})$, which reflects a weighted consideration of semantic similarity (G1, Line 8), available residual capacity (G2, Line 9), and the minimization of truncation penalties (G3, Line 10).
+Through this approach, we aim to ensure that the allocation of documents to sequences is not only computing efficient but also optimally maintains semantic consistency, maximizes capacity utilization, and preserves the integrity of individual documents within each training sequence (Line 11-12).
 The strategy is pivotal in enhancing the overall effectiveness of context sequences for LLMs, by fostering a training environment that is both informationally rich and contextually coherent.
-4. Experiments and Results
-In this section, we describe details of our analytical study of training corpuses (Section 4.1), experimental setup (Section 4.2), the baseline methods we use for comparison (Section 4.3), and experimental results (Section 4.4).
-4.1. Analytical Study on Data Corpus
-The data employed for training originates from a variety of domains, including the web, books, academic papers, code repositories, etc..
-The composition of these sources and their respective weight proportions are illustrated in Figure 3.
-Figure 3. Domain proportion of the training corpus. We have maintained the data distribution ratios across different sources from the previous pretraining stage.
-To provide further insight of our intuition, we undertook a comprehensive analysis, examining both the token quantities within individual documents among diverse domains and assessing the effectiveness of semantic clustering across different methodologies.
-4.1.1. Length Analysis
-We conducted a statistical analysis of the token number characteristics across various domains, as illustrated in Table 1.
-We also categorize the tokenized documents into different length intervals based on token quantities across each domain.
-This process enables us to derive the distribution of token lengths for individual documents situated within specific domains.
-The proportional distributions across domains are presented as a normalized stacked horizontal bar chart in Figure 1, which provides a hierarchical view of the domains according to their prevalence of longer texts.
-It reveals that data originating from the book and academic paper domains predominantly consist of long documents, with more than half of the data exceeding sequence lengths of 64K tokens.
-In contrast, academic papers tend to feature medium-length documents (ranging from 4K to 32K tokens), while web-sourced data are characterized by a preponderance of short documents, almost exclusively comprising texts shorter than 4K tokens.
-Consequently, it is evident that web data does not inherently lend itself to extending model training windows due to its shorter length, necessitating the concatenation of multiple individual documents to form longer sequences suitable for training.
-Domain Book (Chinese) Book (English) Paper (English) Code Paper (Chinese) Web (English) Web (Chinese) Others
-Max Len 24,203,838 8,241,838 10,721,906 1,000,000 1,471,297 1,267,049 80,688 21,839
-Min Len 667 322 127 173 696 289 132 23
-Avg Len 176,156 216,969 20,555 2,578 6,216 886 358 224
+4.3. Semantic Clustering Analysis
 Table 2.
-Token statistics among domains.
-We conducted an analysis of the maximum, minimum, and average token counts from different domains following tokenization.
-Data sourced from the web tends to have a shorter average length, necessitating the concatenation of multiple documents to achieve extended sequences suitable for long-context training.
-The attribute that individual documents originating from books and academic papers tend to have a longer average length significantly increases their propensity to occupy the entirety of a context sequence when training models under expanded contexts, such as 32K or 64K tokens.
-Conversely, web-derived data predominantly exhibits a shorter average document length compared to other domains, necessitating the concatenation of multiple documents to form sequences of adequate length during the training process of extended sequence lengths.
-This observation underscores the intrinsic variability in document length across different domains.
-However, a balanced domain mixture is important for extend context training, where experiments conducted by Fu et al.
-(Fu et al., 2024) demonstrated the importance of upsampling longer sequences while preserving the domain mixture within the pretraining corpora for effective context scaling.
-Therefore, it is necessary to adopt flexible data organization strategies tailored to accommodate the varying length distributions across domains.
-Specifically, for domains characterized by longer documents, such as books and academic papers, leveraging the natural length of these documents can optimize the utilization of context windows during training.
-On the other hand, for web-sourced data, where a majority of the individual documents are insufficient to span the entire context sequence, a strategy involving the concatenation of multiple documents to construct longer sequences becomes imperative.
-This insight holds substantial significance for our work.
-The approach ensures that the training process effectively captures the broader context as well as takes into account the characteristics of different data sources, thereby enhancing the model’s overall ability to understand and generate coherent text across varying lengths and domains.
-Considering the distribution of document lengths and proportion of each domain, we have strategically organized the web-produced English data within our training corpus, which constitutes 30.5% of the total training data.
-4.1.2. Semantic Clustering Analysis
+Document statistics among clusters.
+The clusters formed through ICLM’s greedy traversal of an incomplete graph demonstrate a significant bias in the distribution of document quantities, with 99.8% of clusters containing fewer than 100 documents and 53.56% comprising only a single document.
+In contrast, clusters generated by DataSculpt’s semantic clustering approach display a more balanced distribution of document numbers.
 Method Cluster Number Stats of Docs Number in Clusters
@@
 DataSculpt 611,981 11,747 1 1,827 763 1,479 29
-Table 2.
-Document statistics among clusters.
-The clusters formed through ICLM’s greedy traversal of an incomplete graph demonstrate a significant bias in the distribution of document quantities, with 99.8% of clusters containing fewer than 100 documents and 53.56% comprising only a single document.
-In contrast, clusters generated by DataSculpt’s semantic clustering approach display a more balanced distribution of document numbers.
 (a) T-SNE on clusters built from Algorithm 1.
@@
 Left (4(a)) presents the semantic clustering results generated by the methodology outlined in Algorithm 1, and right (4(b)) illustrates clusters where nodes are united by a shared traversal path, as identified by the greedy graph traversal algorithm discussed in ICLM (Shi et al., 2023).
-In evaluating the effectiveness of semantic clustering, our analysis encompasses two key dimensions: the statistical features related to the document quantities within each cluster, and the application of t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction to project and analyze clusters.
+To demonstrate the robustness of our approach, we compared the effectiveness of semantic clustering across different methodologies.
+Our analysis encompasses two key dimensions: the statistical features related to the document quantities within each cluster, and the application of t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction to project and analyze clusters.
 These clusters were derived from two distinct approaches: the greedy traversal of incomplete graphs as described in ICLM (Shi et al., 2023), for which we refer to the official implementation  [^1] https://github.com/swj0419/in-context-pretraining, and the semantic clustering achieved through an ISODATA variant (Algorithm 1) implemented by our DataSculpt.
-Numerical Narratives.
+4.3.1. Numerical Narratives
 Initially, we performed a quantitative analysis of the number of clusters generated by two distinct methodologies, as detailed in Table 2.
@@
 This strategy efficiently capitalizes on the advantages of clustering similar documents, which significantly enhances the model’s training efficiency by promoting semantic coherence within clusters.
-T-SNE Visualization Journeys.
+4.3.2. T-SNE Visualization Journeys
 Data visualization is crucial for helping decision-making during massive data management (Qin et al., 2019).
-Here we also employed the dimensionality reduction technique t-SNE to visualize both the semantic clusters generated via Algorithm 1 and those identified through greedy graph traversal mechanism on an incomplete graph as described in ICLM.
+Here we employed the dimensionality reduction technique t-SNE to visualize the semantic clusters generated via Algorithm 1 and those identified through greedy graph traversal on an incomplete graph as described in ICLM.
 For the latter approach, each document was connected to its 20 nearest neighbors, facilitating the construction of the incomplete graph.
@@
 The ability of Algorithm 1 to produce semantically concentrated clusters suggests a more robust framework for handling the complexities of large-scale training datasets.
+5. Experiments and Results
+In this section, we describe details of our experimental setup (Section 5.1), the baseline methods we use for comparison (Section 5.2), and experimental results (Section 5.3).
 5.1. Experimental Setup
-We continue pretraining a Transformer decoder-only 7B model (Dong et al., 2024) similar to LLaMA framework (Touvron et al., 2023) by different sources with the same distribution ratios to the previous pretraining stage, where we structured the data derived from the web (English) domain in accordance with a previously defined multi-objective partition allocation strategy (Algorithm 2).
+We continue pretraining a Transformer decoder-only 7B model (Dong et al., 2024) similar to LLaMA framework (Touvron et al., 2023) by different sources with the same distribution ratios to the previous pretraining stage, where we structured the data derived from the web (English) domain in accordance with a previously defined multi-objective partition allocation strategy (Section 4).
 The details of training datasets can be found in Section 4.3.
 Data Processing Details.
-During the data preprocessing phase, we tokenize raw documents from the training corpus (Baichuan, 2023), which collectively encompass 855 billion tokens, and then truncated documents at varying sequence lengths of 16K, 32K, and 64K.
+During the data preprocessing phase, we tokenize raw documents from the training corpus (Baichuan, 2023), which collectively encompass 855 billion tokens, and then truncated the documents at varying sequence lengths of 16K, 32K, and 64K.
 In conducting vector embedding across our training datasets, we employed the BGE-m3 model (Chen et al., 2024b) as our embedding model, encoding each document into a $1024$-dimensional vector.
@@
 Following this, we constructed an HNSW index based on FAISS index (Johnson et al., 2019) to facilitate efficient similarity search for semantic clustering (Section 4.1).
-The greedy semantic-driven partition algorithm (Section 3.4) within each cluster was executed on a 4114 CPU Ray cluster.
+The greedy semantic-driven largest-fit partition algorithm (Section 4.2) within each cluster was executed on a 4114 CPU Ray cluster.
 Model Details.
@@
 We train with 16K, 32K, 64K sequence lengths using 15B tokens for the 7B base model (recall that base model has the 16K context length), applied on 112 H800 GPUs across 14 nodes.
-We employ global batch sizes of 224, 112, and 56 sequences per step for context lengths of 16K, 32K, and 64K, which maintains a uniform number of tokens for model parameter updates.
+We employ global batch sizes of 224, 112, and 56 sequences per step for context lengths of 16K, 32K, and 64K, which aims at maintaining a uniform number of tokens for model parameter updates.
 The learning rate is set to 3e-4 with linear warm-up over the first 1,000 optimization steps and cosine decay schedule.
@@
 General understanding tasks require the model to comprehend the text holistically, extracting key elements and answering queries, which focuses on broad understanding over specific task execution.
-We evaluate the 7B natural language models with zero-shot on MMLU (Hendrycks et al., 2021b, a), TyDiQA (Clark et al., 2020), BBH (Srivastava et al., 2022) and few-shot in-context learning on TREC (Li and Roth, 2002; Hovy et al., 2001).
-MMLU encompasses 57 distinct tasks featuring multiple-choice questions from diverse knowledge domains such as humanities, social sciences and hard sciences.
-TyDiQA is a multilingual question answer dataset containing 204K pairs across 11 topologically diverse languages, which captures language phenomena absent in English-only corpora.
-BBH focuses on a suite of 23 challenging tasks from BIG-Bench that require multi-step reasoning, where previous language model evaluations fell short of average human-rater performance.
-TREC is a text classification challenge aimed at categorizing queries into predefined classes, thereby assessing a model’s ability to accurately classify questions, facilitating the determination of suitable responses or actions.
-Table 3 contains more details about these tasks.
+We evaluate the 7B continual pretrained natural language models with zero-shot on MMLU (Hendrycks et al., 2021), TyDiQA (Clark et al., 2020), BBH (Srivastava et al., 2022) and few-shot in-context learning on TREC (Li and Roth, 2002).
+The evaluation tasks cover different kinds of types, including question answering and text classification.
 Each dataset includes multiple subtasks, and each subtask comes with few-shot examples.
-Dataset # Tasks (Labels) # Instances Answer Type
-MMLU 57 14,079 Letter options
-TyDiQA 9 1,713 Span
-BBH 23 920 COT and answer
-TREC 6(1st)-47(2nd) 500 Letter options
-Table 3. Statistics of evaluation datasets for general understanding. The evaluation tasks cover different kinds of types, including question answering and text classification.
 In Table 3, we compare our models trained via DataSculpt with other baseline models in general understanding across various context window size.
 DataSculpt not only aligns but also slightly outperforms the baselines, particularly in areas requiring multi-hop logical reasoning tasks (+4.88% compared to random sampling, +4.38% compared to ICLM).
+Table 3. Evaluation on general understanding tasks. Despite DataSculpt targeting to long-context utilization, it achieves performance slightly exceeds that of the baselines.
 Context Length Method MMLU TYDIQA BBH TREC
@@
 DataSculpt 46.27 17.55 34.54 71
-Table 4. Evaluation on general understanding tasks. Despite DataSculpt targeting to long-context utilization, it achieves performance slightly exceeds that of the baselines.
-4.4.2. Retrieval Augmentation
-Retrieval augmentation involves LLMs to access external knowledge sources, such as databases or text corpora to retrieve pertinent information as required, effectively extending their memory and knowledge capabilities beyond what is available in their pretrained state.
-In assessing the performance of models on retrieval augmentation tasks, we focus on the models’ synthetic key-value retrieval performance (Liu et al., 2024) under settings of 140 and 300 key-value pairs and conduct a zero-shot evaluation using the TriviaQA dataset (Joshi et al., 2017).
-Further experiments are also undertaken on the “needle in a haystack” task (Kamradt, 2023; Fu et al., 2024) with a maximum of 64K tokens to gain a comprehensive understanding of the effectiveness of these models in identifying specific information within extensive contexts.
-TriviaQA comprises over 650K question-answer-evidence triples and 95K question-answer pairs labeled with evidence passages curated by trivia enthusiasts.
-The key-value retrieval task (Liu et al., 2024) has evolved into a standard diagnostic tool for exploring contextual capabilities, where models are presented with a set of key-value pairs and tasked to retrieve a value corresponding to a specific key from a set of key-value pairs.
-The “needle in a haystack” task is designed to evaluate the models’ capacity for in-context retrieval within extended contexts, further testing their ability to navigate and extract pertinent information from dense information sets.
+Table 4.
+Key-value retrieval performance on dictionaries of 140 ($\approx$ 11.2K tokens) and 300 ($\approx$ 24K tokens) key-value pairs.
+$p_{i}$ represents the scenario where the pertinent information is located at the $i$-th position.
+“-” indicates instances where the result is not applicable, as the tokenized context length exceeds the model’s capacity.
+Each position underwent evaluation across 500 examples.
+The model trained with DataSculpt performs much higher accuracy on hard-to-retrieve positions.
 Context Length Method 140 Key-Value Pairs 300 Key-Value Pairs
@@
 DataSculpt 99.8 96.1 92.4 87.7 61.5 85.5 97.4 64.4 56.8 38 24.8 30.2 72.6 54.89
-Table 5.
-Key-value retrieval performance on dictionaries of 140 ($\approx$ 11.2K tokens) and 300 ($\approx$ 24K tokens) key-value pairs.
-$p_{i}$ represents the scenario where the pertinent information is located at the $i$-th position.
-“-” indicates instances where the result is not applicable, as the tokenized context length exceeds the model’s capacity.
-Each position underwent evaluation across 500 examples.
-The model trained with DataSculpt performs much higher accuracy on hard-to-retrieve positions.
+Table 5. Average performance on retrieval augmentation tasks. DataSculpt outperforms the baselines on almost all datasets, with the mean improvement at 19.33%.
 Context Length Method K-V Retrieval TriviaQA Needle.
@@
 DataSculpt 70.20 86.84 90
-Table 6. Average performance on retrieval augmentation tasks. DataSculpt outperforms the baselines on almost all datasets, with the mean improvement at 19.33%.
 Figure 5.
@@
 The model’s performance in reciting information from the “needle” across various document lengths and positions is assessed on a scale from 1 to 10, represented through a color gradient transitioning from red (score of 1) to green (score of 10).
+5.3.2. Retrieval Augmentation
+Retrieval augmentation involves LLMs to access external knowledge sources, such as databases or text corpora to retrieve pertinent information as required, effectively extending their memory and knowledge capabilities beyond what is available in their pretrained state.
+In assessing the performance of models on retrieval augmentation tasks, we focus on the models’ synthetic key-value retrieval performance (Liu et al., 2024) under settings of 140 and 300 key-value pairs and conduct a zero-shot evaluation using the TriviaQA dataset (Joshi et al., 2017).
+Further experiments are also undertaken on the “needle in a haystack” task (Kamradt, 2023; Fu et al., 2024) with a maximum of 64K tokens to gain a comprehensive understanding of the effectiveness of these models in identifying specific information within extensive contexts.
 Results in Table 4 shows the superior performance of DataSculpt in key-value retrieval tasks across both 140 and 300 key-value pairs configurations, showcasing improvements of up to 59.09% over random sampling and 34.33% over ICLM.
@@
 5.3.3. Summarization
-Summarization tasks involve generation of concise and coherent abstracts from a given source text, which evaluate a model’s capacity to distill key information, maintain the original text’s intent, and present it in a brief and comprehensible manner.
-Compared to traditional QA and text classification tasks, which frequently rely on local context for solutions, summarization requires a broader understanding of the entire document (Bai et al., 2023).
-This capability is more closely associated with LLMs’ long-context capabilities.
-We assess the post-trained models in a zero-shot setting on GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), and employ few-shot in-context learning on SAMSum (Gliwa et al., 2019).
-GovReport is a large-scale dataset for long document summarization, featuring approximately 19.5K U.S.
-government reports paired with expert-crafted abstractive summaries.
-QMSum is a benchmark for query-based multi-domain meeting summarization, encompassing annotated query-summary pairs spanning 232 meetings in diverse domains including product development, academia, and committee discussions.
-SAMSum is specifically curated for dialogue summarization, comprising messenger-style conversations accompanied by human-annotated summaries.
-Additional details about these tasks can be found in Table 7.
+Summarization tasks involve generation of concise and coherent abstracts from a given source text, which serves as a measure of a model’s ability to distill key information, maintain the original text’s intent, and present it in a brief and comprehensible manner.
+Compared to traditional QA and text classification tasks, which predominantly depend on local context for deriving solutions, summarization requires a broader understanding of the entire document (Bai et al., 2023).
+This capability is more closely associated with the long-context processing capabilities of LLMs.
+We assess the continued pretrained models in a zero-shot setting on real-world benchmarks including GovReport (Huang et al., 2021) as well as QMSum (Zhong et al., 2021), and employ few-shot in-context learning on SAMSum (Gliwa et al., 2019).
 Here we utilize the ROUGE-L (Lin, 2004) metric, an N-gram based measure extensively adopted in QA and summarization tasks.
 Improvements in summarization tasks achieved through DataSculpt are documented in Table 6, showcasing enhancements of 21.21% over random sampling and 18.18% over ICLM.
-Dataset Source Avg Len Query Metric
-GovReport Government report 8,734 $\times$ Rouge-L
-QMSum Meeting 10,614
-SAMSum Dialogue 6,258 $\times$
-Table 7.
-Statistics of evaluation datasets for summarization.
-“Source” refers to the provenance of the dataset, and “Avg Len” is calculated based on word counts.
-“Query” signifies whether the dataset conforms to the query-based format.
+Table 6. Evaluation on summarization tasks. DataSculpt consistently surpasses baseline models across nearly all three datasets and various context lengths.
 Context Length Method GovReport QMSum SAMSum
@@
 DataSculpt 29.42 20.25 18.41
-Table 8. Evaluation on summarization tasks. DataSculpt consistently surpasses baseline models across nearly all three datasets and various context lengths.
 5.3.4. Reading Comprehension
-Dataset Avg Len Classification Metric
-NarrativeQA 18,409 Single-Doc F1
-Qasper 3,619
-HotPotQA 9,151 Multi-Doc
-2WikiMultihopQA 4,887
-Table 9. Statistics of evaluation datasets for reading comprehension. “Avg Len” is calculated based on word counts.
+Reading comprehension assesses the capacity of models to generate responses to queries, grounded in the context provided within a given textual passage.
+We evaluate the models’ ability in reading comprehension with zero-shot on NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), HotPotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020).
+NarrativeQA and Qasper fall into the category of single-document question-answering (QA) datasets, whereas HotPotQA and 2WikiMultihopQA are recognized as Wikipedia-based, multi-hop QA datasets.
+Evaluation across these datasets is conducted using the F1 metric, and enhancements attributed to our DataSculpt in reading comprehension tasks are further detailed in Table 7 and Table 8.
+Table 7. In zero-shot evaluations on the validation subsets of NarrativeQA and Qasper, DataSculpt records improvements of 27.06% over random sampling and 13.22% over ICLM.
 Context Length Method NarrativeQA Qasper
@@
 DataSculpt 18.45 9.07
-Table 10. In zero-shot evaluations on the validation subsets of NarrativeQA and Qasper, DataSculpt records improvements of 27.06% over random sampling and 13.22% over ICLM.
+Table 8. Zero-shot F1 performance on the reading comprehension tasks of HotPotQA and 2WikiMultihopQA datasets, with an average improvement of 10.54% on HotPotQA and 6.86% on 2WikiMultihopQA.
 Context Length Method HotPotQA 2WikiMultihopQA
@@
 DataSculpt 9.92 10.23
-Table 11. Zero-shot F1 performance on the reading comprehension tasks of HotPotQA and 2WikiMultihopQA datasets, with an average improvement of 10.54% on HotPotQA and 6.86% on 2WikiMultihopQA.
 Figure 6.
@@
 Among all evaluated training lengths, models developed using the DataSculpt approach surpass baseline-trained models in performance metrics across a diverse range of benchmark categories.
-Reading comprehension assesses the capacity of models to generate responses to queries, grounded in the context provided within a given textual passage.
-We evaluate the models’ ability in reading comprehension with zero-shot on NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), HotPotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020).
-NarrativeQA consists of long stories in the form of books and movie scripts, accompanied by human-crafted questions designed to assess reading comprehension.
-Qasper is a dataset comprising 5,049 questions derived from 1,585 NLP papers.
-The questions are formulated by NLP experts based solely on paper titles and abstracts, while answers are addressed by a different group of NLP practitioners.
-HotpotQA involves 113,000 two-hop question-answer pairs with four key features, authored directly by native speakers given two related paragraphs.
-2WikiMultihopQA is a multi-hop dataset that has comprehensive explanations from question to answer derived from Wikipedia and Wikidata, synthesized with manually designed templates to prevent shortcut-based solutions.
-NarrativeQA and Qasper fall into the category of single-document question-answering (QA) datasets, whereas HotPotQA and 2WikiMultihopQA are recognized as Wikipedia-based, multi-hop QA datasets.
-Evaluation across these datasets is conducted using the F1 metric, as outlined in Table 9.
-Enhancements attributed to DataSculpt in reading comprehension tasks are detailed in Table 10 and Table 11.
 5.3.5. Code Completion
@@
 For code completion tasks, we assess the models using zero-shot learning on the HumanEval dataset (Chen et al., 2021), LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023), where the latter two datasets are specifically designed for long code modeling evaluation.
-HumanEval contains 164 handwritten programming problems, each accompanied by a function signature, a documentation string, the main body of the code, and an average of 7.7 unit tests per challenge to assess correctness.
-LCC is developed by selecting single-file code snippets from GitHub based on length criteria.
-RepoBench-P is curated from GitHub repositories, integrating retrieval and code completion tasks into a cohesive pipeline, where the goal is to first identify and retrieve the most pertinent code snippets based on the context within a file, and then predict the most appropriate subsequent line of code based on the retrieved snippets as well as the in-file context.
-Further information regarding these tasks is provided in Table 12, where we applied execution-based accuracy (Pass@1) for HumanEval evaluation and Edit Sim (Levenshtein distance) for the remaining two code generation tasks.
-Dataset Language Level Metric
-HumanEval Python function Pass@1
-LCC Python/Java/C# file Edit Sim
-RepoBench-P Python/Java repository
-Table 12. Statistics of datasets for code completion. “Language” refers to the programming languages included within the dataset, and “Level” denotes the task’s focus granularity.
+Here we applied execution-based accuracy (Pass@1) for HumanEval evaluation and Edit Sim (Levenshtein distance) for the remaining two code generation tasks.
+Table 9. Performance on code completion, with an enhancement of 4.23% over random sampling and 3.46% on ICLM.
 Context Length Method HumanEval LCC RepoBench-P
@@
 DataSculpt 13.01 64.37 60.04
-Table 13. Performance on code completion, with an enhancement of 4.23% over random sampling and 3.46% on ICLM.
 Results of zero-shot evaluation are presented in Table 9, which reveals a modest advantage of DataSculpt in synergizing average scores for models’ code capabilities.
 While the data was not explicitly structured for the coding domain, advancements in long-context processing achieved through DataSculpt have correspondingly manifested in enhanced coding performance.
+6. Related Work
+6.1. Data-Centric Long-Context Training
+Training Large Language Models (LLMs) with expanded context windows facilitates the execution of tasks that surpass the capabilities defined by the existing paradigm, such as multi-document question answering (Caciularu et al., 2023), long document summarization (Bai et al., 2023), and enhanced retrieval functionalities (Liu et al., 2024).
+However, LLMs’ efficacy in handling long context does not consistently meet the expected standards.
+Data-focused approaches hold substantial importance (Wang et al., 2023) for enhancing LLMs’ long-context capabilities.
+The construction of long-context data involves more than a mere concatenation of short documents, which can lead to cross-document attention within a sequence.
+In these circumstances, formulating high-quality or domain-specific long sequences holds significant research value (de Vries, 2023).
+There have been data-centric efforts made in improving LLMs’ capabilities for long context processing.
+ICLM (Shi et al., 2023) and SPLICE (Staniszewski et al., 2023) combine similar training instances through a sophisticated retrieval process, which emphasizes the critical role of high-quality long-context training data in enhancing the context expansion capabilities of LLMs without necessitating extensive data resources.
+Additionally, Krell et al.
+(Krell et al., 2021) deliberate on the challenges associated with the conventional concat-and-chunk approach, proposing an approximate bin-packing algorithm as a potential solution.
+As an improvement, dataset decomposition (Pouransari et al., 2024) and Best-fit Packing (Ding et al., 2024) both propose the sequence length aware optimization technique to minimize truncations, thereby facilitating more efficient pretraining on long sequences and achieving scalability relative to the size of dataset.
+Moreover, ProLong (Chen et al., 2024a) introduces an evaluation framework for assessing the quality of long-context data, further enriching the field’s methodological toolkit.
+Nonetheless, these methods have not performed a systematic framework tailored for data construction in long context training.
+Our work conceptualizes data management for extended-context training as a multi-objective combinatorial optimization problem, and provide a comprehensive framework for long sequence data organization.
+6.2. MOCO-Augmented LLMs
+Existing works have integrated Multi-objective combinatorial optimization (MOCO) problems with relevant concerns in the LLMs field.
+Wang et al.
+(Wang et al., 2024) employed the MOCO approach within the Reinforcement Learning from Human Feedback (RLHF) framework for LLMs, where the reward models aggregate multiple reward objectives with coefficients into a scalar score.
+Akiba et al.
+(Akiba et al., 2024) viewed model merging as a MOCO challenge, and introduced an innovative approach that leverages evolutionary algorithms to facilitate the automated development of robust foundation models.
+Furthermore, integration of MOCO with data-centric LLMs have been explored for training data selection (Du et al., 2023; Xu et al., 2023; Chen et al., 2024a), which applies heuristic techniques to assign weights to different data evaluation metrics and facilitates a phased resolution approach.
+Considering the extensive scale of training dataset, data construction for long context is treated as a MOCO challenge, for which the coarse-to-fine algorithm is identified as a suitable resolution strategy.
 7. Conclusion
@@
 A pressing challenge lies in organizing long-sequence data for domains lacking extensive text, while ensuring diversity in domain knowledge.
-In response, we introduce DataSculpt, a framework for organizing long-context training data, conceptualizing the task as a multi-objective combinatorial optimization (MOCO) problem.
-This framework considers factors including relevance, homogeneity, integrity, and computational costs associated with data construction, employing a coarse-to-fine method for phased greedy solutions.
-Experimental results indicate that data organized via DataSculpt significantly enhances long-context capabilities without compromising general inference abilities, marking a substantial improvement over existing data management methods.
+In this work, we introduced DataSculpt, a framework for organizing long-context training data, conceptualizing the task as a multi-objective combinatorial optimization problem.
+This framework considered factors including relevance, homogeneity, integrity, and computational costs associated with data construction, employing a coarse-to-fine method for phased greedy solutions.
+Experimental results indicated that data organized via DataSculpt significantly enhances long-context capabilities without compromising general inference abilities, marking a substantial improvement over existing data management methods.